### v2

# On characterization and existence of constrained correlated equilibria in Markov games

Abstract
Markov games with coupling constraints provide a natural framework to study constrained decision-making involving self-interested agents, where the feasibility of an individual agent’s strategy depends on the joint strategies of the others. Such games arise in numerous real-world applications involving safety requirements and budget caps, for example, in environmental management, electricity markets, and transportation systems. While correlated equilibria have emerged as an important solution concept in unconstrained settings due to their computational tractability and amenability to learning, their constrained counterparts remain less explored. In this paper, we study constrained correlated equilibria—feasible policies where any unilateral modifications are either unprofitable or infeasible. We first characterize the constrained correlated equilibrium showing that different sets of modifications result in an equivalent notion, a result which may enable efficient learning algorithms. We then address existence conditions. In particular, we show that a strong Slater-type condition is necessary in games with playerwise coupling constraints, but can be significantly weakened when all players share common coupling constraints. Under this relaxed condition, we prove the existence of a constrained correlated equilibrium.
1 Introduction
Many real-world systems involve multiple self-interested agents that interact with each other in a dynamic environment. Such multi-agent systems can be framed as Markov games, which extend normal-form games to dynamic settings. In many such systems, however, agents are subject to constraints, where the feasibility of an individual agent’s strategy depends on the joint strategies of others. Prominent examples arise in environmental management [33], where countries must jointly ensure that greenhouse emissions are below some threshold, in electricity markets [38], where transmission capacity constraints must be satisfied, or in transportation systems [34], where vehicles must avoid collisions with other vehicles. Such multi-agent systems can be framed as Markov games with coupling constraints [22]. To analyze constrained multi-agent decision making, it is crucial to address constrained equilibrium notions.
The generalized Nash equilibrium was first introduced in [36] for normal-games as a natural extension of its unconstrained counterpart, the Nash equilibrium [35]. [^1] A generalized Nash equilibrium is also referred to as a constrained Nash equilibrium. Several works have since emerged to study the generalized Nash equilibrium, including its characterization and existence in various game settings, ranging from constrained normal-form games [36, 28, 26, 30, 13, 37] to constrained Markov games [3, 4, 24, 25, 40]. However, computing a generalized Nash equilibrium is intractable [21], a property it inherits from the Nash equilibrium. To circumvent this issue in the unconstrained setting, past works have turned to the weaker correlated equilibrium [6, 7], which can be computed efficiently [16, 14, 39]. This motivates studying the constrained correlated equilibrium, which was recently introduced by [17, 11, 8, 31] and to date, has received less attention.
In constrained Markov games, a Markovian constrained correlated equilibrium is a feasible policy such that any unilateral modification is either unprofitable or leads to an infeasible outcome. Different classes of modifications—deterministic vs. stochastic in normal-form games, and additionally, Markovian vs. non-Markovian in Markov games—lead to distinct equilibrium notions. Among these, non-Markovian stochastic modifications form the most general class and define the strongest notion of equilibrium. Understanding whether restricted classes of modifications—such as deterministic or Markovian—can yield equivalent notions of constrained correlated equilibria is a fundamental question. This has practical significance for both computation and learning [19, 17], as limiting the modification set can enable efficient algorithms [10, 27]. For instance, in unconstrained settings, restricting to deterministic modifications enables efficient no-swap regret learning algorithms [10]. However, in the presence of constraints, such restrictions to deterministic modifications have been shown to yield a weaker equilibrium notion [17]. As the structure and characterization of constrained correlated equilibria in terms of possible modifications remain only partially understood, we are led to the following question:
Which classes of modification yield equivalent notions of constrained correlated equilibria?
Apart from characterization, an important question deals with conditions on the existence of constrained correlated equilibria. In general, in constrained games, agents have to satisfy their own constraints, which may depend on other agents’ policies, a setting known as playerwise coupling constraints. For Markov games with playerwise coupling constraints, [3] shows the existence of a constrained Nash equilibrium under a strong Slater’s condition. The existence of a constrained Nash equilibrium implies the existence of a constrained correlated equilibrium. The strong Slater’s condition demands that every agent can modify its policy so that any joint policy is strictly feasible. This is often unrealistic, for example, in transportation systems, it implies that each vehicle can adjust its policy to guarantee collision avoidance with other vehicles.
A possible path to circumvent this issue is to restrict to a setting with common coupling constraints, which are a special case of playerwise coupling constraints, where all agents share the same constraints. Such settings arise in many applications, such as budget caps and safety requirements in environmental management, power grids, and road networks. For normal-form games with common coupling constraints, [11] proves the existence of a constrained correlated equilibrium assuming a jointly feasible policy exists, which is a much weaker condition than the strong Slater’s condition. However, they only consider deterministic modifications and, as discussed above, this leads to a weaker notion of a constrained correlated equilibrium. Thus, it remains unclear whether the relaxation of the strong Slater’s condition is due to the use of common coupling constraints, the restriction to normal-form games, or the limitation to deterministic modifications. Motivated by the above, the second question we address in this paper is the following:
In Markov games with common coupling constraints, does a constrained correlated equilibrium exist without the strong Slater’s condition?
Contributions
In this paper, we consider constrained finite-horizon Markov games with finite state and action spaces and provide answers to the two above questions for this setting. We summarize our contributions as follows:
  1. 1.
    We consider constrained correlated equilibria defined with respect to the most general set of non-Markovian stochastic modifications [17]. We show that it is sufficient to restrict to (i) Markovian stochastic modifications or (ii) convex combinations of Markovian deterministic modifications, thereby characterizing equivalent notions of Markovian constrained correlated equilibria (Theorem 3.3).
  2. 2.
    For Markov games with playerwise coupling constraints, we show that the strong Slater’s condition is necessary for the existence of constrained correlated equilibria (Example 1).
  3. 3.
    For Markov games with common coupling constraints, we prove the existence of a constrained correlated equilibrium under a significantly weakened Slater-type condition (Theorem 4.3). This, together with the second point, shows that the strong Slater’s condition is an artifact of having playerwise coupling constraints.
Notation
Let $\mathbb{N}$ and $\mathbb{R}$ denote the sets of natural and real numbers, respectively. For any $m\in\mathbb{N}$, we define $m]:=\{1,\dots,m\}$. We denote by $x_{m:n}$ the sequence $\{x_{m},x_{m+1},\dots,x_{n}\}$. Let $\mathbf{1}_{n}$ denote the all-ones vector of dimension $n$. Given a finite set $\mathcal{X}$, we denote the indicator function by $\mathds{1}_{\mathcal{X}}(x)$. Furthermore, the probability simplex over $\mathcal{X}$ is denoted by $\Delta({\mathcal{X}})$ and the cardinality of $\mathcal{X}$ is denoted by $|\mathcal{X}|$. Given any two finite sets $\mathcal{X}$ and $\mathcal{Y}$ in $2^{\mathbb{R}^{d}}$, we define the Minkowski sum of the sets $\mathcal{X}$ and $\mathcal{Y}$ as $\mathcal{X}+\mathcal{Y}=\{x+y\mid x\in\mathcal{X},\,y\in\mathcal{Y}\}$. Given a point-to-set mapping $f:\mathcal{X}\to 2^{\mathcal{Y}}$, we say that $f$ is upper semi-compact if it is upper semi-continuous and $f(x)$ is compact for all $x\in\mathcal{X}$ [5, Definition 2.2].
Organization
Section 2 defines a finite-horizon constrained Markov game. Section 3 introduces the constrained correlated equilibrium, where we present equivalent characterizations. Section 4 studies existence conditions, showing that strong Slater’s condition is necessary for playerwise coupling constraints but can be relaxed for common coupling constraints. Section 5 proves our main existence result under the relaxed condition.
2 A finite-horizon Markov game with coupling constraints
A finite-horizon Markov game with coupling constraints is given by the following tuple:
$\{\mathcal{N},H,\mathcal{S},\{\mathcal{A}^{i}\}_{i\in\mathcal{N}},\{r^{i}\}_{i\in\mathcal{N}},\{g^{i,j}\}_{i\in\mathcal{N},j\in[J]},\{c^{i,j}\}_{i\in\mathcal{N},j\in[J]},P,\rho\},$ ($\mathcal{G}$)
where $\mathcal{N}=[N]$ denotes the set of players and $H\in\mathbb{N}$ denotes a finite horizon. The set $\mathcal{S}$ denotes a finite state space and each player $i\in\mathcal{N}$ has a finite action space $\mathcal{A}^{i}$. Let $a=(a^{1},\ldots,a^{N})\in\mathcal{A}$ denote the joint action profile, where $\mathcal{A}=\Pi_{i=1}^{N}A^{i}$ is the joint action space. Similarly, let $a^{-i}=(a^{1},\ldots,a^{i-1},a^{i+1},\ldots,a^{N})\in\mathcal{A}^{-i}=\Pi_{j\neq i}\mathcal{A}^{j}$ denote the joint action profile of all players except player $i$. Furthermore, each player $i\in\mathcal{N}$ has a reward function $r^{i}:=\{r^{i}_{t}\}_{t\in[H]}$, where $r^{i}_{t}:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$. Additionally, each player has $J$ constraint functions $g^{i,j}:=\{g^{i,j}_{t}\}_{t\in[H]}$ for $j\in[J]$, where $g^{i,j}_{t}:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$. The threshold with respect to each player $i$’s constraints $g^{i,j}$ is denoted by $c^{i,j}\in\mathbb{R}$. We refer to the set of constraints $\{g^{i,j}\}_{j\in[J]}$ as playerwise coupling constraints since each player has its own set of constraints that depend on the joint action profile. If, however, the constraint functions and threshold values are equal across all players, i.e., $g^{i,j}=g^{j}$ and $c^{i,j}=c^{j}$ for all $j\in[J]$ and $i\in\mathcal{N}$, then we refer to these types of constraints as common coupling constraints, as each player is subject to the same set of constraints. The transition kernel is denoted by $P:=\{P_{t}\}_{t=1\in[H-1]}$, where $P_{t}:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$. Specifically, $P_{t}(s_{t+1}|s_{t},a_{t})$ denotes the probability of transitioning from state $s_{t}\in\mathcal{S}$ to state $s_{t+1}\in\mathcal{S}$ under the joint action $a_{t}\in\mathcal{A}$. Lastly, $\rho\in\Delta(\mathcal{S})$ denotes the initial state distribution.
For shorthand, we refer to Markov games with coupling constraints as constrained Markov games.
Game dynamics
For $t\in[H]$, let $\pi_{t}:\mathcal{S}\rightarrow\Delta(\mathcal{A})$ denote a mapping from each state $s\in\mathcal{S}$ to a distribution over the joint action space $\mathcal{A}$ at timestep $t$. We define $\pi:=\{\pi_{t}\}_{t\in[H]}$ as a Markovian policy, and denote the set of all such policies by $\Pi_{M}$.
In a constrained Markov game, an initial state $s_{1}$ is drawn from the initial distribution $\rho$, i.e., $s_{1}\sim\rho\in\Delta(\mathcal{S})$. At each timestep $t$, the players select a joint action $a_{t}\sim\pi_{t}(\cdot\lvert s_{t})$ based on policy $\pi\in\Pi_{M}$ and each player receives a reward $r_{t}^{i}(s_{t},a_{t})$ and a set of constraints $\{g_{t}^{i,j}(s_{t},a_{t})\}_{j\in[J]}$. Then, the state $s_{t}$ transitions to a new state $s_{t+1}\sim P_{t}(\cdot\lvert s_{t},a_{t})$. The expected cumulative reward over the horizon $H$ of player $i$ is defined as:
$V^{r^{i}}(\pi):=\mathbb{E}_{s_{1}\sim\rho,a_{t}\sim\pi_{t}(\cdot|s_{t}),s_{t+1}\sim P_{t}(\cdot|s_{t},a_{t})}\left[\sum_{t=1}^{H}r^{i}_{t}(s_{t},a_{t})\right],$
and its expected cumulative constraint value over the horizon $H$ is defined as:
$V^{g^{i,j}}(\pi):=\mathbb{E}_{s_{1}\sim\rho,a_{t}\sim\pi_{t}(\cdot|s_{t}),s_{t+1}\sim P_{t}(\cdot|s_{t},a_{t})}\left[\sum_{t=1}^{H}g^{i,j}_{t}(s_{t},a_{t})\right],\,\forall j\in[J].$
The transition dynamics $P$ and policy $\pi\in\Pi_{M}$ induce a distribution over state-action pairs, the so-called state-action occupancy measure, given by:
$d_{1}^{\pi}(s_{1},a_{1})$ $=\rho(s_{1})\pi_{1}(a_{1}\lvert s_{1}),\,\forall(s_{1},a_{1})\in\mathcal{S}\times\mathcal{A},$
$d_{t}^{\pi}(s_{t},a_{t})$ $=\!\!\!\!\!\!\!\!\!\!\!\!\sum_{(s_{t-1},a_{t-1})\in\mathcal{S}\times\mathcal{A}}\!\!\!\!\!\!\!\!\!\!\!\!d_{t-1}^{\pi}(s_{t-1},a_{t-1})P_{t-1}(s_{t}\lvert s_{t-1},a_{t-1})\pi_{t}(a_{t}\lvert s_{t}),\,\forall(s_{t},a_{t})\in\mathcal{S}\times\mathcal{A},t=2,\dots,H.$
The reward function and constraint functions can then equivalently be expressed as:
$\begin{split}&V^{r^{i}}(\pi)=\sum_{t=1}^{H}\sum_{(s_{t},a_{t})\in\mathcal{S}\times\mathcal{A}}d_{t}^{\pi}(s_{t},a_{t})r_{t}^{i}(s_{t},a_{t}),\\ &V^{g^{i,j}}(\pi)=\sum_{t=1}^{H}\sum_{(s_{t},a_{t})\in\mathcal{S}\times\mathcal{A}}d_{t}^{\pi}(s_{t},a_{t})g_{t}^{i,j}(s_{t},a_{t}),\,\forall j\in[J].\end{split}$ (1)
In a constrained Markov game, each player $i$ is subject to constraints of the form:
$V^{g^{i,j}}(\pi)\geq c^{i,j},\,\forall j\in[J].$ (2)
A policy $\pi$ is said to be $i$-feasible if Inequality (2) above holds for every $j\in[J]$. We denote the set of $i$-feasible policies by:
$\mathcal{C}_{\pi}^{i}:=\{\pi\in\Pi_{M}\mid V^{g^{i,j}}(\pi)\geq c^{i,j},\,\forall j\in[J]\}.$ (3)
Furthermore, a policy $\pi$ is called feasible if it is $i$-feasible for all $i\in\mathcal{N}$. We denote the set of feasible policies by:
$\mathcal{C}_{\pi}:=\{\pi\in\Pi_{M}\mid V^{g^{i,j}}(\pi)\geq c^{i,j},\,\forall j\in[J],\forall i\in\mathcal{N}\}.$ (4)
Observe that $\mathcal{C}_{\pi}=\bigcap_{i\in\mathcal{N}}\mathcal{C}_{\pi}^{i}$. For Markov games with common coupling constraints, we note that $\mathcal{C}_{\pi}^{i}$ reduces to $\mathcal{C}_{\pi}$ for all $i\in\mathcal{N}$ since each player has the same set of constraints.
3 Equivalent notions of constrained correlated equilibrium
An important solution concept in Markov games is the correlated equilibrium. The constrained correlated equilibrium generalizes this notion to constrained Markov games. Such an equilibrium is defined as a policy that is robust against a given set of players’ modifications, in the sense that for each player $i$, deviating to a new policy under a set of modifications will either be not profitable or result in an infeasible policy. In the following, we define possible sets of modifications $\Phi^{i}$, distinguishing between non-Markovian and Markovian, as well as between stochastic and deterministic modifications.
  • For player $i$, the set of non-Markovian stochastic modifications is given by:
    $\Phi_{NM}^{i}:=\{\phi^{i}:=\{\phi_{t}^{i}\}_{t=1}^{H},\text{ where }\phi_{t}^{i}:(\mathcal{S}\times\mathcal{A})^{t-1}\times\mathcal{S}\times\mathcal{A}^{i}\to\Delta(\mathcal{A}^{i}),\,\forall t\in[H]\}.$
  • For player $i$, the set of Markovian stochastic modifications is given by:
    $\Phi_{M}^{i}:=\{\phi^{i}:=\{\phi_{t}^{i}\}_{t=1}^{H},\text{ where }\phi_{t}^{i}:\mathcal{S}\times\mathcal{A}^{i}\to\Delta(\mathcal{A}^{i}),\,\forall t\in[H]\}.$
In the above two definitions, if $\Delta(\mathcal{A}^{i})$ is replaced by $\mathcal{A}^{i}$, then we refer to these sets as non-Markovian deterministic denoted by $\Phi_{NM,det}^{i}$ and Markovian deterministic modifications denoted by $\Phi_{M,det}^{i}$, respectively. Note that $\Phi_{M}^{i}\subseteq\Phi_{NM}^{i}$, $\Phi_{NM,det}^{i}\subseteq\Phi_{NM}^{i}$, and $\Phi_{M,det}^{i}\subseteq\Phi_{M}^{i}$. Given $\phi^{i}\in\Phi^{i}_{NM}$, a modified policy $\phi^{i}\circ\pi:=\{\phi_{t}^{i}\circ\pi_{t}\}_{t=1}^{H}$ is a policy modified by $\phi^{i}$, where $(\hat{a}_{t}^{i},a_{t}^{-i})\sim\pi_{t}$ and then $\phi_{t}^{i}$ modifies $\hat{a}_{t}^{i}$ to another $a_{t}^{i}$ at random. Formally, for all $t\in[H]$, all $a_{t}\in\mathcal{A}$, and all $s_{t}\in\mathcal{S}$, it is given by:
$(\phi_{t}^{i}\circ\pi_{t})(a_{t}\lvert s_{1:t},a_{1:t-1})=\sum_{a_{t}^{i}\in\mathcal{A}^{i}}\phi_{t}^{i}(a_{t}^{i}\lvert s_{1:t},a_{1:t-1},\hat{a}_{t}^{i})\pi_{t}((\hat{a}_{t}^{i},a_{t}^{-i})\lvert s_{t}).$
The constrained correlated equilibrium is defined as:
Definition 3.1 (Constrained correlated equilibrium).
A Markovian policy $\pi\in\Pi_{M}$ is a constrained correlated equilibrium if $\pi$ is feasible, namely, $\pi\in\mathcal{C}_{\pi}$, and if for any player $i\in\mathcal{N}$ the following holds:
$V^{r^{i}}(\pi)\geq\max_{\text{$\phi^{i}\in\Phi_{NM}^{i}$ is $i$-feasible}}V^{r^{i}}(\phi^{i}\circ\pi).$ (5)
In the maximization above, $\phi^{i}$ is said to be $i$-feasible if the modified policy $\phi^{i}\circ\pi$ is $i$-feasible, namely, $\phi^{i}\circ\pi$ satisfies Inequality (2) for every $j\in[J]$.
In unconstrained Markov games, it is known that replacing non-Markovian stochastic modifications $\Phi_{NM}^{i}$ with non-Markovian deterministic modifications $\Phi_{NM,det}^{i}$ results in an equivalent definition of correlated equilibria [17, Theorem 1.1]. We further strengthen this statement by showing that replacing $\Phi_{NM}^{i}$ with Markovian deterministic modifications $\Phi_{M,det}^{i}$ is sufficient for the characterization of correlated equilibria. We provide a proof in Appendix A.
Proposition 3.2.
In unconstrained Markov games, policy $\pi\in\Pi_{M}$ is a correlated equilibrium if for any player $i\in\mathcal{N}$ the following holds:
$V^{r^{i}}(\pi)\geq\max_{\phi^{i}\in\Phi^{i}_{M,det}}V^{r^{i}}(\phi^{i}\circ\pi).$
In the presence of constraints, [17, Theorem 1.2] shows that restricting to deterministic modifications will result in a strictly weaker notion of constrained correlated equibriua. This motivates the question of whether we can limit the class of modifications in Definition 3.1 to get equivalent notions of constrained correlated equilibria. Our result below addresses this.
Theorem 3.3.
The following statements are equivalent:
  1. 1.
    Policy $\pi\in\Pi_{M}$ is a constrained correlated equilibrium as per Definition 3.1.
  2. 2.
    Policy $\pi\in\Pi_{M}$ is feasible and for any player $i\in\mathcal{N}$ the following holds:
    $V^{r^{i}}(\pi)\geq\max_{\text{$\phi^{i}\in\Phi_{M}^{i}$ is $i$-feasible}}V^{r^{i}}(\phi^{i}\circ\pi).$
  3. 3.
    Policy $\pi\in\Pi_{M}$ is feasible and for any player $i\in\mathcal{N}$ the following holds:
    $V^{r^{i}}(\pi)\geq\max_{\text{$\alpha\in\Delta(K^{i})$ is $i$-feasible}}\sum_{k=1}^{K^{i}}\alpha_{k}V^{r^{i}}(\phi^{i}(k)\circ\pi),$
    where $K^{i}:=|\Phi_{M,det}^{i}|$ and $\phi^{i}(k)\in\Phi_{M,det}^{i}$, $\forall k\in[K^{i}]$. We say that $\alpha\in\Delta(K^{i})$ is $i$-feasible if $\alpha\in\{\alpha\mid\sum_{k=1}^{K^{i}}\alpha_{k}V^{g^{i,j}}(\phi^{i}(k)\circ\pi)\geq c^{i,j},\,\forall j\in[J]\}$.
We provide a proof in Appendix A and briefly summarize the results here. The equivalences above imply that in Inequality (5) of Definition 3.1, we can equivalently consider either Markovian modifications $\Phi^{i}_{M}$ or convex combinations of Markovian deterministic modifications $\Phi^{i}_{M,\mathrm{det}}$. The former simplifies the verification of constrained correlated equilibria, while the latter plays a crucial role in our existence proof in the next section.
Remark 3.4.
For learning (constrained) correlated equilibria $\pi\in\Pi_{M}$, Proposition 3.2 and Theorem 3.3 imply that it suffices to consider regret notions with respect to $\Phi^{i}_{M}$, contradicting the conjecture in [15] that such modifications yield a weaker equilibrium notion.
Having characterized constrained correlated equilibria, we now turn to their existence.
4 Existence of a constrained correlated equilibrium
While recent work has focused on the characterization and computation of constrained correlated equilibria [17, 8, 9], existence arguments relied on the framework developed by [3]. However, [3] establishes the existence of constrained correlated equilibria in Markov games with playerwise coupling constraints, under the following assumption: [^2] In [3], the authors prove existence of constrained Nash equilibria for infinite-horizon Markov games, which implies the existence of constrained correlated equilibria. We believe their results can be transferred to the finite-horizon setting by adapting their sensitivity analysis accordingly.
Assumption 4.1 (Strong Slater’s condition).
For any player $i\in\mathcal{N}$ and any policy $\pi\in\Pi_{M}$, there exists a modification $\phi^{i}\in\Phi_{M}^{i}$ such that:
$V^{g^{i,j}}(\phi^{i}\circ\pi)>c^{i,j},\,\forall j\in[J].$
The next example illustrates that even in normal form games the above assumption is necessary for the existence of a constrained correlated equilibrium in settings with playerwise coupling constraints.
Example 1.
Consider a 2-player normal-form game, where each player’s action space is $\mathcal{A}^{1}=\mathcal{A}^{2}=\{1,2\}$. Denote the two players as $P1$ and $P2$, respectively. The reward functions and the playerwise coupling constraint functions for both players are defined as follows:
$r^{1}(a^{1},a^{2})=r^{2}(a^{1},a^{2})=\mathds{1}_{(a^{1},a^{2})=(1,2)},$
$g^{1,1}(a^{1},a^{2})=\mathds{1}_{(a^{1},a^{2})=(1,1)},\quad g^{2,1}(a^{1},a^{2})=\mathds{1}_{(a^{1},a^{2})=(1,2)},$
with constraint thresholds $c^{1,1}=\frac{1}{2}$ and $c^{2,1}=\frac{1}{3}$. It is straightforward to verify that the strong Slater’s condition (Assumption 4.1) does not hold, as there exists no feasible modification for either $P1$ or $P2$ when the policy is $\pi((a^{1},a^{2}))=\mathds{1}_{(a^{1},a^{2})=(2,2)}$. Furthermore, no constrained correlated equilibrium exists in this game since for any feasible policy $\pi$, if $P2$ applies its best $i$-feasible modification with $i=2$, the resulting modified policy strictly increases player $2$’s utility, thereby violating the equilibrium condition. A detailed mathematical justification is provided in Appendix B.1.
In this paper, however, we show that strong Slater’s condition can be relaxed when considering common coupling constraints—where $g^{j}=g^{i,j}$ and $c^{j}=c^{i,j}$ for all $i\in\mathcal{N}$ and all $j\in[J]$. Observe that the terms $i$-feasible and feasible defined in Equation (3) are equivalent, and the set of feasible policies $\mathcal{C}_{\pi}$ defined in Equation (4) reduces to:
$\mathcal{C}_{\pi}:=\{\pi\in\Pi_{M}\mid V^{g^{j}}(\pi)\geq c^{j},\,\forall j\in[J]\}.$ (6)
Now, we present a relaxation of Assumption 4.1.
Assumption 4.2.
For any player $i\in\mathcal{N}$ and any policy $\pi\in\Pi_{M}$ on the boundary of $\mathcal{C}_{\pi}$, i.e., $\pi\in\{\pi\in\Pi_{M}\mid\exists j\in[J]\,\,s.t.\,\,V^{g^{j}}(\pi)=c^{j}\}$, there exists an $\alpha\in\Delta(K^{i})$ with positive weights, namely $\alpha_{k}>0$ for all $k\in[K^{i}]$, such that:
$\sum_{k=1}^{K^{i}}\alpha_{k}V^{g^{j}}(\phi^{i}(k)\circ\pi)\geq c^{j},\,\forall j\in[J],$
where $\phi^{i}(k)\in\Phi^{i}_{M,det}$.
Note that the above assumption is indeed weaker than Assumption 4.1. Namely, it can easily be verified that Assumption 4.1 implies Assumption 4.2. The converse, however, is not true as the following example demonstrates. It shows that there exists a Markov game with common coupling constraints that admits a constrained correlated equilibrium, where the above assumption holds but the strong Slater’s condition is violated.
Example 2.
Consider a 2-player normal-form game with the same reward function as in Example 1. Here, we define the common coupling constraints as follows:
$g^{1}(a^{1},a^{2})=\mathds{1}_{(a^{1},a^{2})=(1,1)},\quad g^{2}(a^{1},a^{2})=\mathds{1}_{(a^{1},a^{2})=(1,2)},$
$g^{3}(a^{1},a^{2})=\mathds{1}_{(a^{1},a^{2})=(2,1)},\quad g^{4}(a^{1},a^{2})=\mathds{1}_{(a^{1},a^{2})=(2,2)},$
with constraint thresholds $c^{j}=\frac{1}{4}$ for all $j\in[4]$. In this setting, the feasible policy set $\mathcal{C}_{\pi}$ contains only one policy, namely the uniform distribution $\pi(a)=\frac{1}{4}$ for all $a\in\mathcal{A}$. As the interior of $\mathcal{C}_{\pi}$ is empty, there exists no modification $\phi^{i}$ for all $i\in[2]$, such that $\phi^{i}\circ\pi$ lies in the interior of $\mathcal{C}_{\pi}$. Consequently, the strong Slater’s condition is violated. However, Assumption 4.2 is satisfied and it is straightforward to verify that the unique feasible policy $\pi$ is indeed a constrained correlated equilibrium. We provide a detailed mathematical justification in Appendix B.2.
Next, we establish our existence result under our relaxed assumption.
Theorem 4.3.
Consider a Markov game with common coupling constraints. Let Assumption 4.2 hold. Then, there exists a policy $\pi\in\Pi_{M}$ which is a constrained correlated equilibrium.
Proof outline
Equilibrium existence proofs typically rely on fixed-point theorems [35, 36, 3]. The key challenge in our proof lies in constructing a point-to-set map $f$ that satisfies the conditions of [32, Kakutani’s fixed point theorem] (see Appendix C), namely, convexity and compactness of the domain and upper semi-continuity of $f$. We then demonstrate that such a fixed point corresponds to a constrained correlated equilibrium.
Considering common coupling constraints ensures that the search for a feasible modification can be restricted to $\mathcal{C}_{\pi}$ for every player $i\in\mathcal{N}$, rather than $\mathcal{C}_{\pi}^{i}$ in the case of playerwise coupling constraints. However, in Markov games $\mathcal{C}_{\pi}$ is generally non-convex [23, Lemma 3]. To overcome this, we consider the set of feasible state-action occupancy $\mathcal{C}_{d}$ given by:
$\mathcal{C}_{d}=\{d^{\pi}\in\Delta(\mathcal{S}\times\mathcal{A})\mid\pi\in\mathcal{C}_{\pi}\},$ (7)
which is convex and compact. This allows us to consider the point-to-set map $f:\mathcal{C}_{d}\to 2^{\mathcal{C}_{d}}$ with domain $\mathcal{C}_{d}$, which is crucial for relaxing the strong Slater’s condition (Assumption 4.1). We provide the full proof in the next section.
Comparison to related work
Altman and Shwartz [3] shows the existence of a constrained Nash equilibrium in Markov games with playerwise coupling constraints. Their proof relies on a fixed-point mapping over the entire policy space $\Pi_{M}$. Firstly, we consider the common coupling constraint setting, which allows us to restrict the search for feasible modifications to $\mathcal{C}_{d}$ in Equation (7). Secondly, we focus on the constrained correlated equilibrium which is crucial for the convexity of $\mathcal{C}_{d}$, a property that does not hold for constrained Nash equilibria [18, Theorem 1].
In the special of normal-form games with common coupling constraints, under the assumption that a feasible policy exists, a relaxation of our Assumption 4.2, [11] establishes the existence of a weaker notion of constrained correlated equilibria by considering only the finite set of deterministic modifications. Note that in normal-form games, there is no horizon, so distinctions between non-Markovian and Markovian modifications are meaningless. In contrast, our setting addresses the more complex case of (i) Markov games with (ii) stochastic modifications. This introduces two key challenges: due to (i) the set $\mathcal{C}_{d}$ in Equation (7) no longer reduces to $\mathcal{C}_{\pi}$ in Equation (6), and (ii) requires searching over infinitely many stochastic modifications. To address this challenge, we leverage the equivalent formulation of the constrained correlated equilibria in Theorem 3.3 and construct a linear program to find the best feasible $\alpha\in\Delta(K^{i})$.
5 Proof of Theorem 4.3
We first establish several preliminary results that ensure the conditions of Kakutani’s fixed point theorem are satisfied. Hereafter, we provide a proof of Theorem 4.3.
5.1 Preliminary results
For a given state-action occupancy measure $d\in\mathcal{C}_{d}$, we define the corresponding feasible Markovian policy $\pi$ via the point-to-set mapping $\Gamma(d):\mathcal{C}_{d}\rightarrow 2^{\mathcal{C}_{\pi}}$ as follows: for every timestep $t\in[H]$, state $s\in\mathcal{S}$ and action $a\in\mathcal{A}$,
$\pi_{t}(a|s)=\begin{cases}\frac{d_{t}(s,a)}{\sum_{a}d_{t}(s,a)},&\text{if }\sum_{a}d_{t}(s,a)\neq 0,\\ \text{arbitrary distribution},&\text{otherwise.}\end{cases}$
The next lemma below establishes properties of the set $\mathcal{C}_{d}$ and mapping $\Gamma(d)$.
Lemma 5.1.
The following holds:
  1. 1.
    The set $\mathcal{C}_{d}$ is compact and convex.
  2. 2.
    The mapping $\Gamma(d)$ is upper semi-compact in $d\in\mathcal{C}_{d}$.
Proof Property 1 follows from [1]. For Property 2, for each $d\in\mathcal{C}_{d}$, $\Gamma(d)$ by construction is a union over $t\in[H]$ and $s\in\mathcal{S}$ consisting of either singletons or $\Delta(\mathcal{A})$, which is compact. Upper semi-continuity in $d\in\mathcal{C}_{d}$ follows from [3, Proposition 3.4].
Given player $i$ and policy $\pi\in\mathcal{C}_{\pi}$, recall that by Theorem 3.3 finding the best feasible modification $\phi^{i}\in\Phi_{M}^{i}$ is equivalent to finding a feasible $\alpha\in\Delta(K^{i})$ which can equivalently be formulated as the following linear program:
Linear Program ($LP^{i}(\pi)$)
Find $\alpha\in\Delta(K^{i})$, where $\Phi^{i}_{M,det}=\{\phi^{i}(k)\}_{k=1}^{K^{i}}$, such that:
$\max_{\alpha}\$$\sum_{k=1}^{K^{i}}\alpha_{k}V^{r^{i}}(\phi^{i}(k)\circ\pi)$
s.t. $\sum_{k=1}^{K^{i}}\alpha_{k}V^{g^{j}}(\phi^{i}(k)\circ\pi)\geq c^{j},\,\forall j\in[J],$ ($LP^{i}(\pi)$)
$\sum_{k=1}^{K^{i}}\alpha_{k}=1,$
$\alpha_{k}\geq 0,\,\forall k\in[K^{i}].$
Let $\Omega^{i}(\pi)$ be the set of optimal solutions of $LP^{i}(\pi)$, and let $\Psi^{i}(\pi)$ denote the optimal value of $LP^{i}(\pi)$. These mappings $\Omega^{i}(\pi)$ and $\Psi^{i}(\pi)$ satisfy the following properties:
Lemma 5.2.
Let Assumption 4.2 hold. Then, the following holds:
  1. 1.
    The optimal set $\Omega^{i}(\pi)$ is convex and upper semi-compact in $\pi\in\mathcal{C}_{\pi}$ for any $i\in\mathcal{N}$.
  2. 2.
    The optimal value $\Psi^{i}(\pi)$ is continuous in $\pi\in\mathcal{C}_{\pi}$ for any $i\in\mathcal{N}$.
We provide a proof in Appendix C.1. The two properties demonstrate compactness and (upper semi-)continuity of the mappings $\Omega^{i}(\pi)$ and $\Psi^{i}(\pi)$. To connect the solution set $\Omega^{i}(\pi)$ of $LP^{i}(\pi)$ to the set of feasible state-action occupnacy measures $\mathcal{C}_{d}$, we define a linear transformation which maps any $\pi\in\mathcal{C}_{\pi}$ to $d\in\mathcal{C}_{d}$:
$A^{i}(\pi):=\bigcup_{\alpha\in\Omega^{i}(\pi)}\left\{\sum_{k\in[K^{i}]}\alpha_{k}d_{t}^{\phi^{i}(k)\circ\pi}(s,a)\right\}_{t\in[H],\,(s,a)\in\mathcal{S}\times\mathcal{A}}.$
The following lemma, for which we provide a proof in Appendix C.2, shows that $A^{i}(\pi)$ defines a valid mapping to $\mathcal{C}_{d}$ and is upper semi-compact.
Lemma 5.3.
For any player $i\in\mathcal{N}$ and policy $\pi\in\mathcal{C}_{\pi}$, the point-to-set mapping $A^{i}(\pi):\mathcal{C}_{\pi}\rightarrow 2^{\mathcal{C}_{d}}$ is upper semi-compact.
5.2 Main proof
Using the results from the previous section, we now provide a proof of our main theorem.
Proof [Proof of Theorem 4.3] Consider the point-to-set map $f:\mathcal{C}_{d}\to 2^{\mathcal{C}_{d}}$ given by:
$f(d):=\bigcup_{i\in\mathcal{N}}f^{i}(\Gamma(d)),$
where for each player $f^{i}:\mathcal{C}_{\pi}\to 2^{\mathcal{C}_{d}}$ is a point-to-set map defined as:
$f^{i}(\pi)$ $:=\left(1-\frac{\Psi^{i}(\pi)-V^{{r}^{i}}(\pi)}{2H}\right)d^{\pi}+\frac{\Psi^{i}(\pi)-V^{{r}^{i}}(\pi)}{2H}A^{i}(\pi).$
First, we will show that for all $i\in\mathcal{N}$ the point-to-set map $f^{i}$ is a valid mapping into $2^{\mathcal{C}_{d}}$ and then that $f^{i}$ is upper semi-compact.
$f^{i}$ is a valid mapping: By assumption $\pi\in\mathcal{C}_{\pi}$ and thus by definition $d^{\pi}\in\mathcal{C}_{d}$. By Lemma 5.3, it holds that $A^{i}(\pi)\subseteq\mathcal{C}_{d}$. Furthermore, it follows that for all $\pi\in\mathcal{C}_{\pi}$:
$\Psi^{i}(\pi)-V^{{r}^{i}}(\pi)=\max_{\alpha\in\Delta(K^{i})\text{ is feasible}}\sum_{k=1}^{K^{i}}\alpha_{k}V^{r^{i}}(\phi^{i}(k)\circ\pi)-V^{{r}^{i}}(\pi)\geq 0.$
The above term can be upper-bounded by: $\Psi^{i}(\pi)-V^{{r}^{i}}(\pi)\leq\Psi^{i}(\pi)\leq H$. Therefore:
$1\geq 1-\frac{\Psi^{i}(\pi)-V^{{r}^{i}}(\pi)}{2H}\geq 1-\frac{H}{2H}=\frac{1}{2}.$
By convexity of $\mathcal{C}_{d}$, shown in Lemma 5.1, the convex combination of $d^{\pi}$ and $d^{i}\in A^{i}(\pi)$ belongs to $\mathcal{C}_{d}$. Therefore, $f^{i}$ is a valid mapping into $2^{\mathcal{C}_{d}}$.
$f^{i}$ is upper semi-compact: By Lemmas 5.2 and 5.3, $\Psi^{i}(\pi)$ is a continuous real-valued function and $A^{i}(\pi)$ is upper semi-compact in $\pi\in\mathcal{C}_{\pi}$. Furthermore, $V^{{r}^{i}}(\pi)$ and $d^{\pi}$ are continuous functions since both can be represented as polynomial functions of $\pi$. Since for every $\pi\in\mathcal{C}_{\pi}$, $f^{i}(\pi)$ is a compact set it follows that $f^{i}$ is upper semi-compact for all $\pi\in\mathcal{C}_{\pi}$.
Next, we will show that the point-to-set map $f:\mathcal{C}_{d}\to 2^{\mathcal{C}_{d}}$ is a valid mapping into $2^{\mathcal{C}_{d}}$ and that $f$ is upper semi-compact.
$f$ is valid mapping and upper semi-compact: The mapping $f$ is a valid mapping into $2^{\mathcal{C}_{d}}$ since $\Gamma$ maps $\mathcal{C}_{d}$ to $2^{\mathcal{C}_{\pi}}$ and $f^{i}$ maps $\mathcal{C}_{\pi}$ to $2^{\mathcal{C}_{d}}$. As $\Gamma(d)$ is upper semi-compact for all $d\in\mathcal{C}_{d}$ by Lemma 5.1, and the composition of two upper semi-compact functions remains upper semi-compact [5, Theorem 5.10], it follows that $f^{i}(\Gamma(d))$ is upper semi-compact for all $d\in\mathcal{C}_{d}$ and $i\in\mathcal{N}$. Furthermore, $f(d)$ is upper semi-compact for all $d\in\mathcal{C}_{d}$ as it is the union of $N$ upper semi-compact functions [5, Theorem 5.4].
In the last step, we will show that for the mapping $f$, a fixed point $d^{*}\in f(d^{*})$ exists and that any corresponding policy $\pi^{*}\in\Gamma(d^{*})$ is a constrained correlated equilibrium.
Existence of a constrained correlated equilibrium: Since by Lemma 5.1, $\mathcal{C}_{d}$ is a compact and convex set and $f(d)$ is upper semi-compact, Kakutani’s fixed point theorem guarantees the existence of a fixed point $d^{*}$ of $f$ such that
$d^{*}\in f(d^{*}).$ (8)
Consider any policy $\pi\in\Gamma(d^{*})$. By Equation (8), for any $i\in\mathcal{N}$, we have that:
$d^{\pi}\in\left(1-\frac{\Psi^{i}(\pi)-V^{{r}^{i}}(\pi)}{2H}\right)d^{\pi}+\frac{\Psi^{i}(\pi)-V^{{r}^{i}}(\pi)}{2H}A^{i}(\pi).$ (9)
To prove $\pi$ is a constrained correlated equilibrium, by Theorem 3.3, we have to show that for any player $i$ and any feasible $\alpha$, the following holds:
$V^{r^{i}}(\pi)\geq\max_{\text{$\alpha\in\Delta(K^{i})$ is feasible}}\sum_{k=1}^{K^{i}}\alpha_{k}V^{r^{i}}(\phi^{i}(k)\circ\pi),$
This is equivalent to showing that $\Psi^{i}(\pi)\leq V^{{r}^{i}}(\pi)$, since $\Psi^{i}(\pi)$ is the optimal solution returned by solving $LP^{i}(\pi)$. We prove this by contradiction. Assume that there exists a player $i\in\mathcal{N}$ and a feasible $\alpha\in\Delta(K^{i})$ such that:
$\sum_{k=1}^{K^{i}}\alpha_{k}V^{r^{i}}(\phi^{i}(k)\circ\pi)>V^{r^{i}}(\pi).$
Combining the above inequality with Equation (9), we have $d^{\pi}\in A^{i}(\pi)$ and by construction of $A^{i}(\pi)$ it follows that: $V^{r^{i}}(\pi)=\Psi^{i}(\pi)$, which is a contradiction. We conclude that any policy $\pi\in\Gamma(d^{*})$ is a constrained correlated equilibrium.
Considering common rather than playerwise coupling constraints is crucial in our analysis to establish the existence of a constrained correlated equilibrium under a significantly weakened Slater-type condition. Moreover, we believe that leveraging a similar analysis, our existence result for common coupling constraints can easily be extended in two directions: Markov games with (i) infinite-horizon and (ii) countable state space and compact action sets. The first direction requires generalizing Theorem 3.3 to the infinite-horizon, and the second direction requires generalizing the arguments related to upper-semi continuity (Lemmas 5.1, 5.2, and 5.3) to countable state spaces and compact action sets.
6 Conclusion and future work
In this paper, we studied constrained correlated equilibria in finite-horizon Markov games with coupling constraints. We addressed two key questions: (i) which classes of modifications yield equivalent definitions of constrained correlated equilibria, and (ii) under what conditions such equilibria exist. Our characterization results show that it suffices to consider either Markovian stochastic modifications or convex combinations of Markovian deterministic modifications, both of which result in equivalent definitions. For existence, we proved that the strong Slater’s condition is necessary in the general setting with playerwise coupling constraints, but can be significantly weakened under common coupling constraints.
A challenging future direction is to establish the existence of constrained Nash equilibria under weakened Slater-type condition in games with common coupling constraints. While our analysis relies on the convexity of the feasible set of state-action occupancy measures, this property does not hold for constrained Nash equilibria requiring a novel analysis to prove existence. Another interesting future direction is to address whether common coupling constraints can bypass the computational intractability found in the computation of constrained correlated with playerwise coupling constraints [9].
Appendix A Proofs of Section 3
We briefly outline the structure of this section. In Section A.1, we show that the set of non-Markovian stochastic modifications $\Phi^{i}_{NM}$ is equivalent, in terms of the state-action occupancy measures they induce, to the set of Markovian stochastic modifications $\Phi^{i}_{M}$. The main technical challenge is constructing an appropriate MDP that enables comparisons between these modification sets.
In Section A.2, we establish a second equivalence—between Markovian stochastic modifications $\Phi^{i}_{M}$, in terms of the state-action occupancy measures they induce, and the convex hull of Markovian deterministic modifications $\Phi^{i}_{M,\mathrm{det}}$. This also requires designing another MDP that enables comparisons between these modification sets.
Finally, in Section A.3, we combine these results to prove Proposition 3.2 and Theorem 3.3.
A.1 Equivalence between $\Phi^{i}_{NM}$ and $\Phi^{i}_{M}$
The aim of this section is to show that the set of state-action occupancy measures induced by non-Markovian stochastic modifications $\Phi^{i}_{{NM}}$ is equivalent to the set of state-action occupancy measures induced by Markovian stochastic modifications $\Phi^{i}_{{M}}$.
To this end, in Section A.1.1, we construct an MDP tailored to player $i$ under policy $\pi$. This enables meaningful comparisons between non-Markovian stochastic modifications and their Markovian stochastic counterparts. Then, in Appendix A.1.2, we prove that this MDP is well-posed (Lemma A.1) and show two key properties: (i) in the constructed MDP, there exists a Markovian stochastic modification $\bar{\phi}^{i}\in\Phi^{i}_{M}$, which induces the same marginal state-action occupancy as the non-Markovian stochastic modification $\phi^{i}\in\Phi^{i}_{NM}$ (Lemma A.2) and (ii) there is a one-to-one mapping from the state-action occupancy measure of the constructed MDP to the state-action occupancy measure of the original Markov game $\mathcal{G}$ (Lemma A.3). Finally, leveraging the above results, in Section A.1.3 we prove the aim of this section (Lemma A.5).
A.1.1 Construction of the first MDP
For any player $i\in\mathcal{N}$ and any policy $\pi\in\Pi_{M}$, we construct a finite-horizon MDP defined by the tuple:
$\left\{\{\tilde{\mathcal{S}}_{t}\cup\{b\}\}_{t\in[H]},\mathcal{A}^{i},H,\tilde{P}^{\pi},\tilde{\rho}\right\}$. (MDP(1))
The state space at timestep $t$ is given by $\tilde{\mathcal{S}_{t}}\cup\{b\}$, where $\tilde{\mathcal{S}}_{t}:=(\mathcal{S}\times\mathcal{A})^{t-1}\times(\mathcal{S}\times\mathcal{A}^{i})$ and $b$ is an auxiliary state. Each element of $\tilde{\mathcal{S}}_{t}$ takes the form $\tilde{s}_{t}=(s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i})$, where $(s_{1:t-1},a_{1:t-1})\in(\mathcal{S}\times\mathcal{A})^{t-1}$ and $(s_{t},a_{t}^{i})\in\mathcal{S}\times\mathcal{A}^{i}$. The time horizon $H$ is the same as in the original Markov game $\mathcal{G}$. The action space $\mathcal{A}^{i}$ corresponds to player $i$’s action set in the original Markov game $\mathcal{G}$. The transition kernel is denoted by $\tilde{P}^{\pi}:=\{\tilde{P}_{t}^{\pi}\}_{t\in[H-1]}$, where each $\tilde{P}_{t}^{\pi}:\tilde{\mathcal{S}_{t}}\times\mathcal{A}^{i}\to[0,1]$ is defined as follows:
$\tilde{P}_{t}^{\pi}\left((s^{\prime}_{1:t},a^{\prime}_{1:t},s_{t+1},a_{t+1}^{i})\mid(s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i}),\hat{a}_{t}^{i}\right)$
$\quad=\frac{1}{|\mathcal{A}^{i}|}P_{t}(s_{t+1}\mid s_{t},(\hat{a}_{t}^{i},a_{t}^{-i}))\pi_{t}((a_{t}^{i},a_{t}^{-i})\mid s_{t})\mathds{1}_{a_{t}^{\prime}=(\hat{a}_{t}^{i},a_{t}^{-i}),(s^{\prime}_{1:t},a^{\prime}_{1:t-1})=(s_{1:t},a_{1:t-1})},$ (10)
$\tilde{P}_{t}^{\pi}\left(b\mid(s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i}),\hat{a}_{t}^{i}\right)=1-\sum_{a_{t}^{-i}}\pi_{t}((a_{t}^{i},a_{t}^{-i})\mid s_{t}).$
We define $b$ to be an absorbing state, such that the transition kernel satisfies:
$\tilde{P}_{t}^{\pi}\left(b\mid b\right)=1.$
Furthermore, we define the initial distribution for all $\tilde{s}_{1}\in\tilde{\mathcal{S}}_{1}\cup\{b\}$ as:
$\tilde{\rho}(\tilde{s}_{1})=\begin{cases}\frac{\rho(s_{1})}{|\mathcal{A}^{i}|},\,&\text{if }\tilde{s}_{1}=(s_{1},a_{1}^{i})\in\tilde{\mathcal{S}}_{1},\\ 0,\,&\text{otherwise}.\end{cases}$ (11)
It is straightforward to verify that $\tilde{\rho}\in\Delta(\tilde{\mathcal{S}}_{1}\cup\{b\})$.
A.1.2 Properties of the first MDP
In the lemma below, we show that that the transition kernel $\tilde{P}_{t}^{\pi}$ of MDP(1) defines a valid probability distribution over the state space $\tilde{\mathcal{S}}_{t+1}\cup\{b\}$ for all $t\in[H-1]$, ensuring that MDP(1) is well-posed.
Lemma A.1.
Suppose $\tilde{P}^{\pi}$ is defined as in Equation (A.1.1). Then, $\tilde{P}^{\pi}_{t}$ is a valid probability distribution in $\Delta(\tilde{\mathcal{S}}_{t+1}\cup\{b\})$ for all $t\in[H-1]$.
Proof For any given timestep $t\in[H-1]$, state $\tilde{s}_{t}=\left(s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i}\right)\in\tilde{S}_{t}$, and action $\hat{a}_{t}^{i}\in\mathcal{A}_{i}$, it holds that:
$\sum_{\tilde{s}_{t+1}\in\tilde{\mathcal{S}}_{t+1}\cup\{b\}}\tilde{P}_{t}^{\pi}\left(\tilde{s}_{t+1}|\left(s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i}\right),\hat{a}_{t}^{i}\right)$
$=$ $\sum_{s^{\prime}_{1:t},a^{\prime}_{1:t}}\sum_{s_{t+1},a_{t+1}^{i}}\tilde{P}_{t}^{\pi}\left(\left(s^{\prime}_{1:t},a^{\prime}_{1:t},s_{t+1},a_{t+1}^{i}\right)|\left(s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i}\right),\hat{a}_{t}^{i}\right)$
$+\tilde{P}_{t}^{\pi}\left(b|\left(s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i}\right),\hat{a}_{t}^{i}\right)$
$=$ $\sum_{s^{\prime}_{1:t},a^{\prime}_{1:t}}\sum_{s_{t+1},a_{t+1}^{i}}\frac{1}{|\mathcal{A}^{i}|}P_{t}(s_{t+1}|s_{t},(\hat{a}_{t}^{i},a_{t}^{-i}))\pi_{t}((a_{t}^{i},a_{t}^{-i})|s_{t})\mathds{1}_{a_{t}^{\prime}=(\hat{a}_{t}^{i},a_{t}^{-i}),(s^{\prime}_{1:t},a^{\prime}_{1:t-1})=(s_{1:t},a_{1:t-1})}$
$+1-\sum_{a_{t}^{-i}}\pi_{t}((a_{t}^{i},a_{t}^{-i})|s_{t})$
$=$ $\sum_{a_{t+1}^{i}}\frac{1}{|\mathcal{A}^{i}|}\sum_{a_{t}^{-i}}\sum_{s_{t+1}}P_{t}(s_{t+1}|s_{t},(\hat{a}_{t}^{i},a_{t}^{-i}))\pi_{t}((a_{t}^{i},a_{t}^{-i})|s_{t})+1-\sum_{a_{t}^{-i}}\pi_{t}((a_{t}^{i},a_{t}^{-i})|s_{t})$
$=$ $\sum_{a_{t}^{-i}}\pi_{t}((a_{t}^{i},a_{t}^{-i})|s_{t})+1-\sum_{a_{t}^{-i}}\pi_{t}((a_{t}^{i},a_{t}^{-i})|s_{t})$
$=$ $1.$
Furthermore, for the auxiliary state $b$ and timestep $t\in[H-1]$, it holds that:
$\sum_{\tilde{s}\in\tilde{\mathcal{S}}_{t+1}\cup\{b\}}\tilde{P}_{t}^{\pi}\left(\tilde{s}\mid b\right)=\tilde{P}_{t}^{\pi}(b\mid b)=1.$
We conclude that $\tilde{P}_{t}^{\pi}$ is a valid probability distribution in $\Delta(\tilde{\mathcal{S}}_{t+1}\cup\{b\})$.
We note that any non-Markovian stochastic modification $\phi^{i}\in\Phi^{i}_{NM}$, where $\phi^{i}:=\{\phi_{t}^{i}\}_{t=1}^{H}$ with $\phi_{t}^{i}:(\mathcal{S}\times\mathcal{A})^{t-1}\times\mathcal{S}\times\mathcal{A}^{i}\to\Delta(\mathcal{A}^{i})$, constitutes a valid policy in MDP(1). Specifically, given any $\tilde{s}_{t}\in\tilde{\mathcal{S}}_{t}$ action $\hat{a}_{t}^{i}$ is selected according to $\hat{a}_{t}^{i}\sim\phi_{t}^{i}(\cdot\mid\tilde{s}_{t})$. Furthermore, for any given non-Markovian modification $\phi^{i}\in\Phi_{{NM}}^{i}$, we construct a corresponding Markovian modification $\bar{\phi}^{i}\in\Phi_{{M}}^{i}$, where each $\bar{\phi}_{t}^{i}:\mathcal{S}\times\mathcal{A}^{i}\to\Delta(\mathcal{A}^{i})$ is defined as follows:
$\bar{\phi}_{t}^{i}(\hat{a}_{t}^{i}\mid s_{t},a_{t}^{i})=\begin{cases}\frac{\sum\limits_{s_{1:t-1},a_{1:t-1}}\tilde{d}_{t}^{\phi^{i}}((s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i}),\hat{a}_{t}^{i})}{\sum\limits_{\hat{a}_{t}^{i}}\sum\limits_{s_{1:t-1},a_{1:t-1}}\tilde{d}_{t}^{\phi^{i}}((s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i}),\hat{a}_{t}^{i})},&\text{if the denominator}>0,\\ \text{any distribution in }\Delta(\mathcal{A}^{i}),&\text{otherwise}.\end{cases}$ (12)
We note that $\bar{\phi}^{i}$ is also a valid policy in MDP(1). Namely, for any state $\tilde{s}_{t}=(s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i})$ in $\tilde{\mathcal{S}}_{t}$, we can define the policy as:
$\bar{\phi}^{i}_{t}(\hat{a}_{t}^{i}\mid\tilde{s}_{t}):=\bar{\phi}^{i}_{t}(\hat{a}_{t}^{i}\mid s_{t},a_{t}^{i}).$
In the next lemma, we show that the original modification $\phi^{i}\in\Phi^{i}_{NM}$ and the corresponding Markovian modification $\bar{\phi}^{i}\in\Phi^{i}_{M}$ constructed in (12) induce the same marginal state-action occupancy measure over $(s_{t},a_{t}^{i})$ and $\hat{a}_{t}^{i}$ for every timestep $t$ in MDP(1).
Lemma A.2.
For all $h\in[H]$, $(s_{h},a_{h}^{i})\in\mathcal{S}\times\mathcal{A}^{i}$, and $\hat{a}_{h}^{i}\in\mathcal{A}_{i}$, the following equality holds:
$\sum_{s_{1:h-1},a_{1:h-1}}\!\!\!\tilde{d}_{h}^{\phi^{i}}((s_{1:h-1},a_{1:h-1},s_{h},a_{h}^{i}),\hat{a}_{h}^{i})=\!\!\!\sum_{s_{1:h-1},a_{1:h-1}}\!\!\!\tilde{d}_{h}^{\bar{\phi}^{i}}((s_{1:h-1},a_{1:h-1},s_{h},a_{h}^{i}),\hat{a}_{h}^{i}),$ (13)
where $\tilde{d}_{h}^{\phi^{i}}$ and $\tilde{d}_{h}^{\bar{\phi}^{i}}$ denote the state-action occupancy measures at timestep $h$ of MDP(1), induced by the modifications $\phi^{i}$ and $\bar{\phi}^{i}$, respectively.
Proof We prove the above lemma by induction.
Base case: For $h=1$ note that:
$\sum_{\hat{a}_{1}^{i}}\tilde{d}_{1}^{\phi^{i}}((s_{1},a_{1}^{i}),\hat{a}_{1}^{i})=\sum_{\hat{a}_{1}^{i}}\tilde{\rho}(s_{1},a_{1}^{i})\phi_{1}^{i}(\hat{a}_{1}^{i}|s_{1},a_{1}^{i})=\tilde{\rho}(s_{1},a_{1}^{i}).$ (14)
Then, if $\tilde{\rho}(s_{1},a_{1}^{i})\neq 0$, we can express $\tilde{d}_{1}^{\bar{\phi}^{i}}((s_{1},a_{1}^{i}),\hat{a}_{1}^{i})$ as follows:
$\tilde{d}_{1}^{\bar{\phi}^{i}}((s_{1},a_{1}^{i}),\hat{a}_{1}^{i})$ $=\tilde{\rho}(s_{1},a_{1}^{i})\bar{\phi}_{1}^{i}(\hat{a}^{i}|s_{1},a_{1}^{i})$
$=\tilde{\rho}(s_{1},a_{1}^{i})\frac{\tilde{d}_{1}^{\phi^{i}}((s_{1},a_{1}^{i}),\hat{a}_{1}^{i})}{\sum_{\hat{a}_{1}^{i}}\tilde{d}_{1}^{\phi^{i}}((s_{1},a_{1}^{i}),\hat{a}_{1}^{i})}$
[…]
If $\tilde{\rho}(s_{1},a_{1}^{i})=0$, then we obtain
$\tilde{d}_{1}^{\bar{\phi}^{i}}((s_{1},a_{1}^{i}),\hat{a}_{1}^{i})=\tilde{d}_{1}^{\phi^{i}}((s_{1},a_{1}^{i}),\hat{a}_{1}^{i})=0.$
Induction step: Assume that Equation (13) holds for all $(s_{h},a_{h}^{i})\in\mathcal{S}_{h}\times\mathcal{A}^{i}$, all $\hat{a}_{h}^{i}\in\mathcal{A}_{i}$, and all $h\in[t-1]$. We now show that it also holds for $t$. For $\tilde{d}_{t}^{\phi^{i}}$, we have
$\sum_{\hat{a}_{t}^{i}}\sum_{s_{1:t-1},a_{1:t-1}}\tilde{d}_{t}^{\phi^{i}}((s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i}),\hat{a}_{t}^{i})$
$=$ $\sum_{\hat{a}_{t}^{i}}\sum_{s^{\prime}_{1:t-1},a^{\prime}_{1:t-1}}\sum_{s_{1:t-2},a_{1:t-2}}\sum_{s_{t-1},a_{t-1}^{i}}\sum_{\hat{a}_{t-1}^{i}}\tilde{d}_{t-1}^{\phi^{i}}((s_{1:t-2},a_{1:t-2},s_{t-1},a_{t-1}^{i}),\hat{a}_{t-1}^{i})$
$\times\tilde{P}_{t-1}^{\pi}((s^{\prime}_{1:t-1},a^{\prime}_{1:t-1},s_{t},a_{t}^{i})|(s_{1:t-2},a_{1:t-2},s_{t-1},a_{t-1}^{i}),\hat{a}_{t-1}^{i})\phi^{i}_{t}(\hat{a}_{t}^{i}|(s^{\prime}_{1:t-1},a^{\prime}_{1:t-1},s_{t},a_{t}^{i}))$
$\overset{(i)}{=}$ $\sum_{s^{\prime}_{1:t-1},a^{\prime}_{1:t-1}}\sum_{s_{1:t-2},a_{1:t-2}}\sum_{s_{t-1},a_{t-1}^{i}}\sum_{\hat{a}_{t-1}^{i}}\tilde{d}_{t-1}^{\phi^{i}}((s_{1:t-2},a_{1:t-2},s_{t-1},a_{t-1}^{i}),\hat{a}_{t-1}^{i})$
$\times\tilde{P}_{t-1}^{\pi}((s^{\prime}_{1:t-1},a^{\prime}_{1:t-1},s_{t},a_{t}^{i})|(s_{1:t-2},a_{1:t-2},s_{t-1},a_{t-1}^{i}),\hat{a}_{t-1}^{i})$
$\overset{(ii)}{=}$ $\frac{1}{|\mathcal{A}^{i}|}\sum_{s_{t-1}}\sum_{(\hat{a}_{t-1}^{i},a_{t-1}^{-i})}\sum_{a_{t-1}^{i}}P_{t-1}(s_{t}|s_{t-1},(\hat{a}_{t-1}^{i},a_{t-1}^{-i}))\pi_{t-1}((a_{t-1}^{i},a_{t-1}^{-i})|s_{t-1})$
$\times\sum_{s_{1:t-2},a_{1:t-2}}\tilde{d}_{t-1}^{\phi^{i}}((s_{1:t-2},a_{1:t-2},s_{t-1},a_{t-1}^{i}),\hat{a}_{t-1}^{i}).$ (15)
In step $(i)$, we apply that $\sum_{\hat{a}_{t}^{i}}\phi^{i}_{t}(\hat{a}_{t}^{i}|(s^{\prime}_{1:t-1},a^{\prime}_{1:t-1},s_{t},a_{t}^{i}))=1$ and in step $(ii)$, we use the definition of $\tilde{P}^{\pi}_{t-1}$. For $\tilde{d}_{t}^{\bar{\phi}^{i}}$, we write
$\sum_{s_{1:t-1},a_{1:t-1}}\tilde{d}_{t}^{\bar{\phi}^{i}}((s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i}),\hat{a}_{t}^{i})$
$=$ $\bar{\phi}_{t}^{i}(\hat{a}_{t}^{i}|s_{t},a_{t}^{i})\sum_{s^{\prime}_{1:t-1},a^{\prime}_{1:t-1}}\sum_{s_{1:t-2},a_{1:t-2}}\sum_{s_{t-1},a_{t-1}^{i}}\sum_{\hat{a}_{t-1}^{i}}\tilde{d}_{t-1}^{\bar{\phi}^{i}}((s_{1:t-2},a_{1:t-2},s_{t-1},a_{t-1}^{i}),\hat{a}_{t-1}^{i})$
$\times\tilde{P}_{t-1}^{\pi}((s^{\prime}_{1:t-1},a^{\prime}_{1:t-1},s_{t},a_{t}^{i})|(s_{1:t-2},a_{1:t-2},s_{t-1},a_{t-1}^{i}),\hat{a}_{t-1}^{i})$
$\overset{(i)}{=}$ $\frac{\bar{\phi}_{t}^{i}(\hat{a}_{t}^{i}|s_{t},a_{t}^{i})}{|\mathcal{A}^{i}|}\sum_{s_{t-1}}\sum_{(\hat{a}_{t-1}^{i},a_{t-1}^{-i})}\sum_{a_{t-1}^{i}}P_{t-1}(s_{t}|s_{t-1},(\hat{a}_{t-1}^{i},a_{t-1}^{-i}))\pi_{t-1}((a_{t-1}^{i},a_{t-1}^{-i})|s_{t-1})$
$\times\sum_{s_{1:t-2},a_{1:t-2}}\tilde{d}_{t-1}^{\bar{\phi}^{i}}((s_{1:t-2},a_{1:t-2},s_{t-1},a_{t-1}^{i}),\hat{a}_{t-1}^{i})$
$\overset{(ii)}{=}$ $\frac{\bar{\phi}_{t}^{i}(\hat{a}_{t}^{i}|s_{t},a_{t}^{i})}{|\mathcal{A}^{i}|}\sum_{s_{t-1}}\sum_{(a_{t-1}^{-i},\hat{a}_{t-1}^{i})}\sum_{a_{t-1}^{i}}P_{t-1}(s_{t}|s_{t-1},(\hat{a}_{t-1}^{i},a_{t-1}^{-i}))\pi_{t-1}((a_{t-1}^{i},a_{t-1}^{-i})|s_{t-1})$
$\times\sum_{s_{1:t-2},a_{1:t-2}}\tilde{d}_{t-1}^{\phi^{i}}((s_{1:t-2},a_{1:t-2},s_{t-1},a_{t-1}^{i}),\hat{a}_{t-1}^{i})$
$\overset{(iii)}{=}$ ${\bar{\phi}_{t}^{i}(\hat{a}_{t}^{i}|s_{t},a_{t}^{i})}\left(\sum_{\hat{a}_{t}^{i}}\sum_{s_{1:t-1},a_{1:t-1}}\tilde{d}_{t}^{\phi^{i}}((s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i}),\hat{a}_{t}^{i})\right).$ (16)
In step $(i)$, we use the definition of $\tilde{P}^{\pi}_{t-1}$ and in step $(ii)$, we apply the following equation, which holds by induction:
$\sum_{s_{1:t-2},a_{1:t-2}}\!\!\!\!\tilde{d}_{t-1}^{\bar{\phi}^{i}}((s_{1:t-2},a_{1:t-2},s_{t-1},a_{t-1}^{i}),\hat{a}_{t-1}^{i})=\!\!\!\!\sum_{s_{1:t-2},a_{1:t-2}}\!\!\!\!\tilde{d}_{t-1}^{\phi^{i}}((s_{1:t-2},a_{1:t-2},s_{t-1},a_{t-1}^{i}),\hat{a}_{t-1}^{i}).$
In step $(iii)$, we plug in Equation (15). If Equation (15) is non-zero, we can substitute the definition of $\bar{\phi}^{i}$ from Equation (12) into Equation (16). This results in:
$\sum_{s_{1:t-1},a_{1:t-1}}\tilde{d}_{t}^{\bar{\phi}^{i}}\left((s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i}),\hat{a}_{t}^{i}\right)=\sum_{s_{1:t-1},a_{1:t-1}}\tilde{d}_{t}^{\phi^{i}}\left((s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i}),\hat{a}_{t}^{i}\right).$
Otherwise, if Equation (15) is zero, by Equation (16) we have the following:
$\sum_{s_{1:t-1},a_{1:t-1}}\tilde{d}_{t}^{\bar{\phi}^{i}}\left((s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i}),\hat{a}_{t}^{i}\right)=0.$
Meanwhile, if Equation (15) is zero, we obtain
$\sum_{s_{1:t-1},a_{1:t-1}}\tilde{d}_{t}^{\phi^{i}}\left((s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i}),\hat{a}_{t}^{i}\right)=0,$
which holds as each state-action distribution is non-negative. Therefore, we conclude that:
$\sum_{s_{1:t-1},a_{1:t-1}}\tilde{d}_{t}^{\bar{\phi}^{i}}\left((s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i}),\hat{a}_{t}^{i}\right)=\sum_{s_{1:t-1},a_{1:t-1}}\tilde{d}_{t}^{\phi^{i}}\left((s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i}),\hat{a}_{t}^{i}\right).$
In the lemma below, we demonstrate a one-to-one mapping between the state-action occupancy measure of policy $\phi^{i}$ in MDP(1) and that of the modified policy $\phi^{i}\circ\pi$ in the Markov game $\mathcal{G}$.
Lemma A.3.
For any player $i\in\mathcal{N}$, Markovian policy $\pi\in\Pi_{M}$, and non-Markovian stochastic modification $\phi^{i}\in\Phi_{NM}^{i}$, the state-action occupancy measure of the modified policy $\phi^{i}\circ\pi$ equals:
$\!\!\!\!\!\!d_{h}^{\phi^{i}\circ\pi}({s}_{h},(\hat{a}_{h}^{i},a_{h}^{-i}))=|\mathcal{A}^{i}|^{h}\sum_{a_{h}^{i}}\sum_{s_{1:h-1},a_{1:h-1}}\!\!\!\!\!\!\!\tilde{d}_{h}^{\phi^{i}}((s_{1:h-1},a_{1:h-1},s_{h},a_{h}^{i}),\hat{a}_{h}^{i})\pi_{h}((a_{h}^{i},a_{h}^{-i})|s_{h}),$ (17)
for all $(s_{h},(\hat{a}_{h}^{i},a_{h}^{-i}))\in\mathcal{S}\times\mathcal{A}$ and $h\in[H]$, where ${d}_{h}^{\phi^{i}\circ\pi}$ denotes the state-action occupancy measure at timestep $h$ of Markov game $\mathcal{G}$ and $\tilde{d}_{h}^{\phi^{i}}$ denotes the state-action occupancy measure at timestep $h$ of MDP(1).
Remark A.4.
The lemma above also holds for any Markovian stochastic modification $\bar{\phi}^{i}\in\Phi_{M}^{i}$ since $\Phi_{M}^{i}$ is a subset of $\Phi_{NM}^{i}$ and $\bar{\phi}^{i}$ constitutes a valid policy in MDP(1).
Proof We prove the above lemma by induction.
Base case: For $h=1$, we have that for any $(s_{1},(\hat{a}_{1}^{i},a_{1}^{-i}))\in\mathcal{S}\times\mathcal{A}$ the following holds:
$|\mathcal{A}^{i}|\sum_{a_{1}^{i}}\tilde{d}_{1}^{\phi^{i}}((s_{1},a_{1}^{i}),\hat{a}^{i})\pi_{1}((a_{1}^{i},a_{1}^{-i})\mid s_{1})$
$\overset{(i)}{=}$ $|\mathcal{A}^{i}|\sum_{a_{1}^{i}}\tilde{\rho}(s_{1},a_{1}^{i})\phi^{i}_{1}(\hat{a}^{i}_{1}|s_{1},a_{1}^{i})\pi_{1}((a_{1}^{i},a_{1}^{-i})\mid s_{1})$
$\overset{(ii)}{=}$ $\sum_{a_{1}^{i}}\rho(s_{1})\phi_{1}^{i}(\hat{a}^{i}_{1}|s_{1},a_{1}^{i})\pi_{1}((a_{1}^{i},a_{1}^{-i})\mid s_{1})$
$=$ $d_{1}^{\phi^{i}\circ\pi}({s}_{1},(a_{1}^{-i},\hat{a}_{1}^{i})),$
where in step $(i)$ we express $\tilde{d}_{1}^{\phi^{i}}$ by the dynamics of MDP(1). In step $(ii)$ we plug in the definition of $\tilde{\rho}$ given in Equation (11).
Induction step: Assume that Equation (17) holds for all $(s_{h},(a_{h}^{-i},\hat{a}_{h}^{i}))\in\mathcal{S}\times\mathcal{A}$ and all $h\in[t-1]$, then we show that Equation (17) also holds for $t$. We can rewrite $\tilde{d}_{t}^{\phi^{i}}$ as:
$\tilde{d}_{t}^{\phi^{i}}((s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i}),\hat{a}_{t}^{i})$
$\overset{(i)}{=}$ $\sum_{\bar{s}_{1:t-2},\bar{a}_{1:t-2}}\sum_{\bar{s}_{t-1}}\sum_{\bar{a}_{t-1}^{i}}\sum_{\hat{a}_{t-1}^{i}}\tilde{d}_{t-1}^{\phi^{i}}((\bar{s}_{1:t-2},\bar{a}_{1:t-2},\bar{s}_{t-1},\bar{a}_{t-1}^{i}),\hat{a}_{t-1}^{i})$
$\times\tilde{P}_{t-1}^{\pi}(s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i}|(\bar{s}_{1:t-2},\bar{a}_{1:t-2},\bar{s}_{t-1},\bar{a}_{t-1}^{i}),\hat{a}_{t-1}^{i})\phi^{i}_{t}(\hat{a}_{t}^{i}|s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i})$
$\overset{(ii)}{=}$ $|\mathcal{A}^{i}|^{-1}\sum_{\bar{s}_{1:t-2},\bar{a}_{1:t-2}}\sum_{\bar{s}_{t-1}}\sum_{\bar{a}_{t-1}^{i}}\sum_{\hat{a}_{t-1}^{i}}\tilde{d}_{t-1}^{\phi^{i}}((\bar{s}_{1:t-2},\bar{a}_{1:t-2},\bar{s}_{t-1},\bar{a}_{t-1}^{i}),\hat{a}_{t-1}^{i})\pi_{t-1}((\bar{a}_{t-1}^{i},\bar{a}_{t-1}^{-i})\mid\bar{s}_{t-1})$
$\times P_{t-1}(s_{t}|\bar{s}_{t-1},(\hat{a}_{t-1}^{i},\bar{a}_{t-1}^{-i}))\mathds{1}_{a_{t-1}=(\hat{a}_{t-1}^{i},\bar{a}_{t-1}^{-i}),(s_{1:t-1},a_{1:t-2})=(\bar{s}_{1:t-1},\bar{a}_{1:t-2})}\phi^{i}_{t}(\hat{a}_{t}^{i}|s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i})$
$\overset{(iii)}{=}$ $|\mathcal{A}^{i}|^{-1}\sum_{\bar{a}_{t-1}^{i}}\tilde{d}_{t-1}^{\phi^{i}}((s_{1:t-2},a_{1:t-2},s_{t-1},\bar{a}_{t-1}^{i}),a_{t-1}^{i})\pi_{t-1}((\bar{a}_{t-1}^{i},a_{t-1}^{-i})\mid s_{t-1})P_{t-1}(s_{t}|s_{t-1},a_{t-1})$
$\times\phi^{i}_{t}(\hat{a}_{t}^{i}|s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i}).$
In step $(i)$, we express the state-action distribution using the transition dynamics, in step $(ii)$, we apply the definition of $\tilde{P}^{\pi}_{t-1}$, and in step $(iii)$, we sum only over the indicator function with non-zero value. Using the above equation, we obtain:
$|\mathcal{A}^{i}|^{t}\sum_{a_{t}^{i}}\sum_{s_{1:t-1},a_{1:t-1}}\tilde{d}_{t}^{\phi^{i}}((s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i}),\hat{a}_{t}^{i})\pi_{t}((a_{t}^{i},a_{t}^{-i})|s_{t})$
$=$ $|\mathcal{A}^{i}|^{t-1}\sum_{a_{t}^{i}}\sum_{s_{1:t-1},a_{1:t-1}}\sum_{\bar{a}_{t-1}^{i}}\tilde{d}_{t-1}^{\phi^{i}}((s_{1:t-2},a_{1:t-2},s_{t-1},\bar{a}_{t-1}^{i}),a_{t-1}^{i})\pi_{t-1}((\bar{a}_{t-1}^{i},a_{t-1}^{-i})\mid s_{t-1})$
$\times P_{t-1}(s_{t}|s_{t-1},a_{t-1})\phi^{i}_{t}(\hat{a}_{t}^{i}|s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i})\pi_{t}((a_{t}^{i},a_{t}^{-i})|s_{t})$
$\overset{(i)}{=}$ $\sum_{a_{t}^{i}}\sum_{s_{1:t-1},a_{1:t-1}}d_{t-1}^{\phi^{i}\circ\pi}(s_{1:t-1},a_{1:t-1})P_{t-1}(s_{t}|s_{t-1},a_{t-1})\phi^{i}_{t}(\hat{a}_{t}^{i}|s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i})\pi_{t}((a_{t}^{i},a_{t}^{-i})|s_{t})$
$=$ $\sum_{s_{1:t-1},a_{1:t-1}}d_{t-1}^{\phi^{i}\circ\pi}(s_{1:t-1},a_{1:t-1})P_{t-1}(s_{t}|s_{t-1},a_{t-1})\sum_{a_{t}^{i}}\phi^{i}_{t}(\hat{a}_{t}^{i}|s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i})\pi_{t}((a_{t}^{i},a_{t}^{-i})|s_{t})$
$=$ $\sum_{s_{1:t-1},a_{1:t-1}}d_{t}^{\phi^{i}\circ\pi}(s_{1:t-1},a_{1:t-1},{s}_{t},(\hat{a}_{t}^{i},a_{t}^{-i}))$
$=$ $d_{t}^{\phi^{i}\circ\pi}({s}_{t},(\hat{a}_{t}^{i},a_{t}^{-i})),$
where in step $(i)$, we apply Equation (17) which holds by induction:
$|\mathcal{A}^{i}|^{t-1}\sum_{\bar{a}_{t-1}^{i}}\sum_{s_{1:t-2},a_{1:t-2}}\tilde{d}_{t-1}^{\phi^{i}}((s_{1:t-2},a_{1:t-2},s_{t-1},\bar{a}_{t-1}^{i}),a_{t-1}^{i})\pi_{t-1}((\bar{a}_{t-1}^{i},a_{t-1}^{-i})|s_{t-1})$
$=$ $d_{t-1}^{\phi^{i}\circ\pi}({s}_{t-1},({a}_{t-1}^{i},a_{t-1}^{-i}))$
$=$ $\sum_{s_{1:t-2},a_{1:t-2}}d_{t-1}^{\phi^{i}\circ\pi}(s_{1:t-1},a_{1:t-1}).$
A.1.3 State-action occupancy measure of $\Phi^{i}_{NM}$ and $\Phi^{i}_{M}$
In this section, we utilize the results established in Appendix A.1.2 to prove that the set of modifications $\Phi^{i}_{\mathrm{NM}}$ and $\Phi^{i}_{\mathrm{M}}$ induce the same state-action occupancy measure over the Markov game $\mathcal{G}$.
Lemma A.5.
Given player $i\in\mathcal{N}$ and policy $\pi\in\Pi_{M}$, for any non-Markovian modification $\phi^{i}\in\Phi_{NM}^{i}$, there exists a corresponding Markovian modification $\bar{\phi}^{i}\in\Phi_{M}^{i}$ such that:
$d_{t}^{\phi^{i}\circ\pi}(s_{t},a_{t})=d_{t}^{\bar{\phi}^{i}\circ\pi}(s_{t},a_{t}),\quad\forall(s_{t},a_{t})\in\mathcal{S}\times\mathcal{A},\forall t\in[H].$
Proof Given player $i$ and policy $\pi$, any non-Markovian stochastic modification $\phi^{i}\in\Phi_{NM}^{i}$ in the original Markov game $\mathcal{G}$ constitutes a valid policy in MDP(1), since each $\phi_{t}^{i}$ maps from $\tilde{\mathcal{S}}_{t}$ to $\Delta(\mathcal{A}^{i})$.
In Lemma A.3, we established the following equality between the distribution over the state-action space in the original Markov game $\mathcal{G}$ and MDP(1): for all $(s_{t},(\hat{a}_{t}^{i},a_{t}^{-i}))\in\mathcal{S}\times\mathcal{A}$ and all $t\in[H]$ it holds that:
$d_{t}^{\phi^{i}\circ\pi}({s}_{t},(\hat{a}_{t}^{i},a_{t}^{-i}))=|\mathcal{A}^{i}|^{t}\sum_{a_{t}^{i}}\sum_{s_{1:t-1},a_{1:t-1}}\!\!\!\!\!\tilde{d}_{t}^{\phi^{i}}((s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i}),\hat{a}_{t}^{i})\pi_{t}((a_{t}^{i},a_{t}^{-i})|s_{t}).$ (18)
Recall Equation (12), where, for any given modification $\phi^{i}$, we constructed $\bar{\phi}^{i}$ which corresponds to a Markovian stochastic modification in the original Markov game $\mathcal{G}$. Then, similarly by Lemma A.3, for all $(s_{t},(\hat{a}_{t}^{i},a_{t}^{-i}))\in\mathcal{S}\times\mathcal{A}$ and all $t\in[H]$ it holds that:
$d_{t}^{\bar{\phi}^{i}\circ\pi}({s}_{t},(\hat{a}_{t}^{i},a_{t}^{-i}))=|\mathcal{A}^{i}|^{t}\sum_{a_{t}^{i}}\sum_{s_{1:t-1},a_{1:t-1}}\!\!\!\!\!\tilde{d}_{t}^{\bar{\phi}^{i}}((s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i}),\hat{a}_{t}^{i})\pi_{t}((a_{t}^{i},a_{t}^{-i})|s_{t}).$ (19)
Furthermore, by Lemma A.2, the following equality holds for all $t\in[H]$, $(s_{t},a_{t}^{i})\in\mathcal{S}\times\mathcal{A}^{i}$, and $\hat{a}_{t}^{i}\in\mathcal{A}_{i}$:
$\sum_{s_{1:t-1},a_{1:t-1}}\tilde{d}_{t}^{\phi^{i}}((s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i}),\hat{a}_{t}^{i})=\sum_{s_{1:t-1},a_{1:t-1}}\tilde{d}_{t}^{\bar{\phi}^{i}}((s_{1:t-1},a_{1:t-1},s_{h},a_{t}^{i}),\hat{a}_{t}^{i}).$ (20)
Combining Equations (18), (19), and (20), it holds that:
$d_{t}^{\bar{\phi}^{i}\circ\pi}({s}_{t},(\hat{a}_{t}^{i},a_{t}^{-i}))=$ $|\mathcal{A}^{i}|^{t}\sum_{a_{t}^{i}}\sum_{s_{1:t-1},a_{1:t-1}}\tilde{d}_{t}^{\bar{\phi}^{i}}((s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i}),\hat{a}_{t}^{i})\pi_{t}((a_{t}^{i},a_{t}^{-i})|s_{t})$
$=$ $|\mathcal{A}^{i}|^{t}\sum_{a_{t}^{i}}\sum_{s_{1:t-1},a_{1:t-1}}\tilde{d}_{t}^{\phi^{i}}((s_{1:t-1},a_{1:t-1},s_{t},a_{t}^{i}),\hat{a}_{t}^{i})\pi_{t}((a_{t}^{i},a_{t}^{-i})|s_{t})$
$=$ $d_{t}^{\phi^{i}\circ\pi}({s}_{t},(\hat{a}_{t}^{i},a_{t}^{-i})).$
A.2 Equivalence between $\Phi^{i}_{M}$ and convex combination of $\Phi^{i}_{M,{det}}$
The aim of this section is to show that the set of state-action occupancy measures induced by Markovian stochastic modifications $\Phi^{i}_{\mathrm{M}}$ is equivalent to the convex hull of state-action occupancy measures induced by Markovian deterministic modifications $\Phi^{i}_{\mathrm{M,det}}$.
The proof parallels the arguments in Appendix A.1. In Appendix A.2.1, we begin by constructing an MDP tailored to player $i$ under policy $\pi$, enabling direct comparisons between Markovian stochastic and Markovian deterministic modifications. We show that this MDP is well-defined (Lemma A.6) and admits a one-to-one correspondence between its state-action occupancy measure and that of the original Markov game (Lemma A.7). Finally, in Section A.2.2 we prove the aim of this section by leveraging the classical result of MDPs, namely that stochastic Markov policies are equivalent to convex combinations of deterministic ones in terms of occupancy measures (Lemma A.8).
A.2.1 Construction and properties of the second MDP
Similarly to MDP(1), for any player $i\in\mathcal{N}$ and any policy $\pi\in\Pi_{M}$, we construct a finite-horizon MDP defined by the tuple:
$\left\{\bar{\mathcal{S}}\cup\{b\},\mathcal{A}^{i},H,\bar{P}^{\pi},\bar{\rho}\right\}.$ (MDP(2))
The state space is given by $\bar{\mathcal{S}}\cup\{b\}$, where $\bar{\mathcal{S}}:=\mathcal{S}\times\mathcal{A}^{i}$ and $b$ is an auxiliary state. Each element of $\bar{\mathcal{S}}$ takes the form $\bar{s}=(s,a^{i})$. The time horizon $H$ is the same as in the original Markov game $\mathcal{G}$. The action space $\mathcal{A}^{i}$ corresponds to player $i$’s action set in the original Markov game $\mathcal{G}$. The transition kernel is denoted by $\bar{P}^{\pi}:=\{\bar{P}_{t}^{\pi}\}_{t\in[H-1]}$, where each $\bar{P}_{t}^{\pi}:\bar{\mathcal{S}}\times\mathcal{A}^{i}\to[0,1]$ is defined as follows:
$\bar{P}_{t}^{\pi}\left((s_{t+1},a_{t+1}^{i})\mid(s_{t},a_{t}^{i}),\hat{a}_{t}^{i}\right)=\frac{1}{|\mathcal{A}^{i}|}\sum_{a_{t}^{-i}}P_{t}(s_{t+1}\mid s_{t},(\hat{a}_{t}^{i},a_{t}^{-i}))\pi_{t}((a_{t}^{i},a_{t}^{-i})\mid s_{t}),$ (21)
$\bar{P}_{t}^{\pi}\left(b\mid(s_{t},a_{t}^{i}),\hat{a}_{t}^{i}\right)=1-\sum_{a_{t}^{-i}}\pi_{t}((a_{t}^{i},a_{t}^{-i})\mid s_{t}).$
We define $b$ to be an absorbing state, such that the transition kernel satisfies:
$\bar{P}_{t}^{\pi}\left(b\mid b\right)=1.$
Furthermore, we define the initial distribution for all $\bar{s}_{1}\in\bar{\mathcal{S}}\cup\{b\}$ as:
$\bar{\rho}(\bar{s}_{1})=\begin{cases}\frac{\rho(s_{1})}{|\mathcal{A}^{i}|},\,&\text{if }\bar{s}_{1}=(s_{1},a_{1}^{i})\in\bar{\mathcal{S}}_{1},\\ 0,\,&\text{otherwise}.\end{cases}$
It is straightforward to verify that $\bar{\rho}\in\Delta(\bar{\mathcal{S}}\cup\{b\})$. Lemma A.6 ensures that MDP(2) is well-posed.
Lemma A.6.
Suppose $\bar{P}^{\pi}$ is defined as in Equation (21). Then, $\bar{P}^{\pi}_{t}$ is a valid probability distribution in $\Delta(\bar{\mathcal{S}}\cup\{b\})$ for all $t\in[H-1]$.
The proof follows from the same arguments as in the proof of Lemma A.1.
We note that any Markovian stochastic modification $\bar{\phi}^{i}\in\Phi^{i}_{M}$, i.e., $\bar{\phi}^{i}:=\{\bar{\phi}_{t}^{i}\}_{t=1}^{H}$ with $\bar{\phi}_{t}^{i}:\mathcal{S}\times\mathcal{A}^{i}\to\Delta(\mathcal{A}^{i})$, constitutes a valid policy in MDP(2). In the following lemma, we demonstrate a one-to-one mapping between the state-action occupancy measure of the modified policy $\bar{\phi}^{i}\circ\pi$ in the original Markov game $\mathcal{G}$ and that of the MDP(2).
Lemma A.7.
For any player $i\in\mathcal{N}$, Markovian policy $\pi\in\Pi_{M}$, and Markovian stochastic modification $\bar{\phi}^{i}\in\Phi_{M}^{i}$, the state-action occupancy measure of the modified policy $\bar{\phi}^{i}\circ\pi$ equals:
$\!\!\!\!\!\!d_{h}^{\bar{\phi}^{i}\circ\pi}({s}_{h},(\hat{a}_{h}^{i},a_{h}^{-i}))=|\mathcal{A}^{i}|^{h}\sum_{a_{h}^{i}}\bar{d}_{h}^{\bar{\phi}^{i}}((s_{h},a_{h}^{i}),\hat{a}_{h}^{i})\pi_{h}((a_{h}^{i},a_{h}^{-i})|s_{h}),$
for all $(s_{h},(\hat{a}_{h}^{i},a_{h}^{-i}))\in\mathcal{S}\times\mathcal{A}$ and $h\in[H]$, where $\bar{d}_{h}^{\bar{\phi}^{i}}(\tilde{s},a^{i})$ denotes the state-action occupancy measure at timestep $h$ of MDP(2).
The proof follows from the same arguments as in the proof of Lemma A.3.
A.2.2 State-action occupancy measure of $\Phi^{i}_{M}$ and $\Phi^{i}_{M,det}$
Here, we utilize the results of the previous section to establish a connection between any Markovian stochastic modification $\bar{\phi}^{i}\in\Phi_{M}^{i}$ and the set of deterministic modifications $\Phi_{M,det}^{i}$ at the level of state-action occupancy of the Markov game $\mathcal{G}$.
Lemma A.8.
Given player $i$ and policy $\pi\in\Pi_{M}$, for any Markovian modification $\bar{\phi}^{i}\in\Phi^{i}_{M}$, there exists a vector $\alpha\in\Delta(K^{i})$ such that:
$d_{t}^{\bar{\phi}^{i}\circ\pi}(s_{t},a_{t})=\sum_{k=1}^{K^{i}}\alpha_{k}d_{t}^{\bar{\phi}^{i}(k)\circ\pi}(s_{t},a_{t}),\,\forall(s_{t},a_{t})\in\mathcal{S}\times\mathcal{A},\forall t\in[H],$
where $\bar{\phi}^{i}(k)\in\Phi_{M,det}^{i}$, $k\in[K^{i}]$. The converse of this statement also holds.
Proof Given player $i$ and policy $\pi\in\Pi_{M}$, consider any modification $\bar{\phi}^{i}\in\Phi_{M}^{i}$. Such a modification of the Markov game $\mathcal{G}$ defines a valid policy for MDP(2).
In the MDP setting, we leverage that Markovian stochastic policies and convex combinations of Markovian deterministic policies are equivalent in terms of the state-action occupancy measures they induce (see [29, Theorem 5.2]). Namely, there exists a vector $\alpha\in\Delta(K^{i})$ such that:
$\bar{d}_{t}^{\bar{\phi}^{i}}((s_{t},a_{t}^{i}),\hat{a}_{t}^{i})=\sum_{k=1}^{K^{i}}\alpha_{k}\,\bar{d}_{t}^{\bar{\phi}^{i}(k)}((s_{t},a_{t}^{i}),\hat{a}_{t}^{i}),$ (22)
for all $(s_{t},a_{t}^{i})\in\bar{\mathcal{S}}$, all $\hat{a}_{t}^{i}\in\mathcal{A}^{i}$, and all $t\in[H]$, where each $\bar{\phi}^{i}(k)\in\Phi^{i}_{M,det}$ is a Markovian deterministic policy of MDP(2). Note that since Equation (22) holds with equality, the reverse statement is also true. Namely, given a vector $\alpha\in\Delta(K^{i})$, there exists a $\bar{\phi}^{i}\in\Phi_{M}^{i}$ such that Equation (22) holds.
Then, by Lemma A.7, the corresponding state-action occupancy measure of the modified policy $\bar{\phi}^{i}\circ\pi$ which includes $\bar{\phi}^{i}(k)\circ\pi$ in the original Markov game $\mathcal{G}$ is given by:
$d_{t}^{\bar{\phi}^{i}\circ\pi}(s_{t},(\hat{a}_{t}^{i},a_{t}^{-i}))=|\mathcal{A}^{i}|^{t}\sum_{a_{t}^{i}}\sum_{s_{1:t-1},a_{1:t-1}}\bar{d}_{t}^{\bar{\phi}^{i}}((s_{t},a_{t}^{i}),\hat{a}_{t}^{i})\,\pi_{t}((a_{t}^{i},a_{t}^{-i})\mid s_{t}).$
for all $s_{t}\in\mathcal{S}$, all $(\hat{a}_{t}^{i},a_{t}^{-i})\in\mathcal{A}$, and all $t\in[H]$. Combining the above two expressions, we conclude the proof.
A.3 Main proofs of Section 3
Leveraging the results in Section A.1 and A.2, we first provide the main proof of Proposition 3.2 and then we provide the main proof of Theorem 3.3.
Now, we are ready to prove Proposition 3.2.
Proof [Proof for Proposition 3.2] Given player $i$ and policy $\pi\in\Pi_{M}$, we note that any Markovian stochastic modification $\bar{\phi}^{i}\in\Phi^{i}_{M}$ constitutes a valid policy in MDP(2). For this MDP, we define a reward function $\bar{r}^{i,\pi}$ as follows: for $t\in[H]$,
$\bar{r}_{t}^{i,\pi}(\bar{s}_{t},\hat{a}_{t}^{i})=|\mathcal{A}^{i}|^{t}\sum_{a_{t}^{-i}}r_{t}^{i}(s_{t},(\hat{a}_{t}^{i},a_{t}^{-i}))\pi_{t}((a_{t}^{i},a_{t}^{-i})\mid s_{t}),\,\,\forall\bar{s}_{t}=(s_{t},a_{t}^{i})\in\bar{\mathcal{S}},$
and $\bar{r}_{t}^{i,\pi}(b)=0$. The expected cumulative reward $V^{\bar{r}^{i,\pi}}(\bar{\phi}^{i})$ of MDP(2) is given by:
$V^{\bar{r}^{i,\pi}}(\bar{\phi}^{i})$
$\overset{(i)}{=}$ $\sum_{t=1}^{H}\sum_{s_{t},a_{t}^{i}}\sum_{\hat{a}_{t}^{i}}\bar{d}_{t}^{\bar{\phi}^{i}}((s_{t},a_{t}^{i}),\hat{a}_{t}^{i})\bar{r}_{t}^{i,\pi}((s_{t},a_{t}^{i}),\hat{a}_{t}^{i})$
$\overset{(ii)}{=}$ $\sum_{t=1}^{H}|\mathcal{A}^{i}|^{t}\sum_{s_{t},a_{t}^{i}}\sum_{\hat{a}_{t}^{i}}\bar{d}_{t}^{\bar{\phi}^{i}}((s_{t},a_{t}^{i}),\hat{a}_{t}^{i})\left(\sum_{a_{t}^{-i}}r_{t}^{i}(s_{t},(\hat{a}_{t}^{i},a_{t}^{-i}))\pi_{t}((a_{t}^{i},a_{t}^{-i})\mid s_{t})\right)$
${=}$ $\sum_{t=1}^{H}\sum_{s_{t}}\sum_{(\hat{a}_{t}^{i},a_{t}^{-i})}r_{t}^{i}(s_{t},(\hat{a}_{t}^{i},a_{t}^{-i}))\left(|\mathcal{A}^{i}|^{t}\sum_{a_{t}^{i}}\bar{d}_{t}^{\bar{\phi}^{i}}((s_{t},a_{t}^{i}),\hat{a}_{t}^{i})\pi_{t}((a_{t}^{i},a_{t}^{-i})|s_{t})\right)$
$\overset{(iii)}{=}$ $\sum_{t=1}^{H}\sum_{s_{t}}\sum_{(\hat{a}_{t}^{i},a_{t}^{-i})}r_{t}^{i}(s_{t},(\hat{a}_{t}^{i},a_{t}^{-i}))\left(d_{t}^{\bar{\phi}^{i}\circ\pi}({s}_{t},(\hat{a}_{t}^{i},a_{t}^{-i}))\right)$
$=$ $V^{r^{i}}(\bar{\phi}^{i}\circ\pi),$
where $(i)$ holds since the reward $\bar{r}_{t}^{i,\pi}(b)$ is zero for all $t\in[H]$. Furthermore, in step $(ii)$, we substitute the definition of $\bar{r}_{t}^{i,\pi}$ and in step $(iii)$, we apply Lemma A.7. The above equation implies that:
$\max_{\bar{\phi}^{i}\in\Phi^{i}_{M}}V^{r^{i}}(\bar{\phi}^{i}\circ\pi)=\max_{\bar{\phi}^{i}\in\Phi^{i}_{M}}V^{\bar{r}^{i,\pi}}(\bar{\phi}^{i}).$ (23)
Furthermore, in MDPs there always exists an optimal Markovian deterministic policy which maximizes the expected cumulative reward [1] such that:
$\max_{\bar{\phi}^{i}\in\Phi^{i}_{M}}V^{\bar{r}^{i,\pi}}(\bar{\phi}^{i})=\max_{\bar{\phi}^{i}(k)\in\Phi^{i}_{M,det}}V^{\bar{r}^{i,\pi}}(\bar{\phi}^{i}(k)).$ (24)
Combining Equations (23) and (24), we conclude the proof.
Next, we provide the main proof of Theorem 3.3.
Proof [Proof of Theorem 3.3]
  • Equivalence between the first and second statement: By Lemma A.5, for any $\phi^{i}\in\Phi^{i}_{NM}$ there exists a $\bar{\phi}^{i}\in\Phi_{M}$ such that:
    $d_{t}^{\phi^{i}\circ\pi}(s_{t},a_{t})=d_{t}^{\bar{\phi}^{i}\circ\pi}(s_{t},a_{t}),\quad\forall(s_{t},a_{t})\in\mathcal{S}_{t}\times\mathcal{A},\forall t\in[H].$
    Since the reward $V^{r^{i}}(\phi^{i}\circ\pi)$ and the constraint functions $V^{g^{i,j}}(\phi^{i}\circ\pi)$ are linear functions of the state-action occupancy measure as seen in Equation (1), we have:
    $V^{r^{i}}(\phi^{i}\circ\pi)=V^{r^{i}}(\bar{\phi}^{i}(k)\circ\pi)\,\,\text{and}\,\,V^{g^{i,j}}(\phi^{i}\circ\pi)=V^{g^{i,j}}(\bar{\phi}^{i}(k)\circ\pi),\,\forall j\in[J].$
    This implies the equivalence between the first and the second statement in Theorem 3.3.
  • Equivalence between the first and third statement: By Lemma A.8, for any $\bar{\phi}^{i}\in\Phi^{i}_{M}$, there exists a vector $\alpha\in\Delta(K^{i})$ such that:
    $d_{t}^{\bar{\phi}^{i}\circ\pi}(s_{t},a_{t})=\sum_{k=1}^{K^{i}}\alpha_{k}d_{t}^{\bar{\phi}^{i}(k)\circ\pi}(s_{t},a_{t}),\,\forall(s_{t},a_{t})\in\mathcal{S}\times\mathcal{A},\forall t\in[H],$
    where $\bar{\phi}^{i}(k)\in\Phi_{M,det}^{i}$. By the same linearity arguments as above, we conclude the equivalence between the first and third statements in Theorem 3.3.
Appendix B Verification of examples
In the following, we provide detailed mathematical justifications for Examples 1 and 2. Since we focus on the normal-form game setting (i.e., with the time horizon $H=1$ and one state), we omit the time index $t$ and the state $s$ for simplicity in this section.
B.1 Verification of Example 1
Proof Consider a 2-player normal-form game, where each player’s action space is $\mathcal{A}^{1}=\mathcal{A}^{2}=\{1,2\}$. Denote by $P1$ and $P2$ the first and the second player, respectively. The reward and the playerwise coupling constraint functions of each player are given by:
$P1$ $P2$ 11 22
11 (0,0) (1,1)
22 (0,0) (0,0)
Table 1: Rewards $r^{1}$ and $r^{2}$.
$P1$ $P2$ 11 22
11 (1,0) (0,1)
22 (0,0) (0,0)
Table 2: Constraints $g^{1,1}$ and $g^{2,1}$.
We consider policies $\pi\in\Delta(\mathcal{A})$ of the form $\pi((1,1),(1,2),(2,1),(2,2))=(x,y,z,1-x-y-z)$, represented as:
$P1$ $P2$ 11 22
11    x y
22 z 1-x-y-z
P1 faces the following optimization problem:
$\max_{\pi\in\Delta(\mathcal{A})}V^{r^{1}}(\pi),\quad\text{s.t. }V^{g^{1,1}}(\pi)\geq\frac{1}{2}.$
Plugging in the values of reward and constraint functions, this is equivalent to:
$\max_{\pi\in\Delta(\mathcal{A})}y,\quad\text{s.t. }x\geq\frac{1}{2}.$
P2 faces the following optimization problem:
$\max_{\pi\in\Delta(\mathcal{A})}V^{r^{2}}(\pi),\quad\text{s.t. }V^{g^{2,1}}(\pi)\geq\frac{1}{3}.$
Plugging in the values of reward and constraint functions, this is equivalent to:
$\max_{\pi\in\Delta(\mathcal{A})}y,\quad\text{s.t. }y\geq\frac{1}{3}.$
We now establish the following claims:
  1. 1.
    The strong Slater’s condition does not hold: Consider the policy $\pi=(0,0,0,1)$, then for any modification $\phi^{1}\in\Phi^{1}_{NM}$ of P1, the modified policy is of the form $\phi^{1}\circ\pi=(0,y,0,1-y)\in\Delta(\mathcal{A})$, which is $i$-infeasible for P1, $i=1$. Similarly, we can verify that P2 does not have an $i$-feasible modification, $i=2$, either.
  2. 2.
    A constrained correlated equilibrium does not exist: By Definition 3.1, a constrained correlated equilibrium does not exist if for any feasible policy, there exists a player $i$ with an $i$-feasible modification that strictly increases player $i$’s utility. Given any feasible policy $\pi$, player 2’s best modification is
    $\phi^{2}(\hat{a}^{2}|a^{2})=\mathbf{1}_{\{\hat{a}^{2}=2\}},$
    namely, always modifying to action $a^{2}=2$. Such modification results in the policy of the form:
    $P1$ $P2$ 11 22
    11   0 y+x
    22 0 1-x-y
    This modification is always $2$-feasible and strictly increases player 2’s reward.
    Therefore, for every feasible policy, player 2 admits a $2$-feasible modification that improves its reward. Hence, no constrained correlated equilibrium exists. However, this new policy $\phi^{2}\circ\pi$ is not $i$-feasible where $i=1$ since P1’s constraint is violated.
B.2 Verification of Example 2
Proof Consider a 2-player normal-form game with the same reward function as in Example 1. And we consider four common coupling constraint functions given by:
$g^{1}(a^{1},a^{2})=\mathds{1}_{(a^{1},a^{2})=(1,1)},\quad g^{2}(a^{1},a^{2})=\mathds{1}_{(a^{1},a^{2})=(1,2)},$
$g^{3}(a^{1},a^{2})=\mathds{1}_{(a^{1},a^{2})=(2,1)},\quad g^{4}(a^{1},a^{2})=\mathds{1}_{(a^{1},a^{2})=(2,2)}.$
We consider policies $\pi\in\Delta(\mathcal{A})$ of the form:
$\pi=\pi((1,1),(1,2),(2,1),(2,2))=(x,y,z,1-x-y-z).$
P1 faces the following optimization problem:
$\max_{\pi\in\Delta(\mathcal{A})}V^{r^{1}}(\pi),\quad\text{s.t. }V^{g^{j}}(\pi)\geq\frac{1}{4},\,\forall j\in[4].$
Plugging in the values of reward and constraint functions, this is equivalent to:
$\max_{\pi\in\Delta(\mathcal{A})}y,\quad\text{s.t. }x\geq\frac{1}{4},\,y\geq\frac{1}{4},\,z\geq\frac{1}{4},\,1-x-y-z\geq\frac{1}{4}.$
Analogously, P2 faces the following optimization problem:
$\max_{\pi\in\Delta(\mathcal{A})}V^{r^{2}}(\pi),\quad\text{s.t. }V^{g^{j}}(\pi)\geq\frac{1}{4},\,\forall j\in[4].$
Plugging in the values of reward and constraint functions, this is equivalent to:
$\max_{\pi\in\Delta(\mathcal{A})}y,\quad\text{s.t. }x\geq\frac{1}{4},\,y\geq\frac{1}{4},\,z\geq\frac{1}{4},\,1-x-y-z\geq\frac{1}{4}.$
We now establish the following claims:
  1. 1.
    The first condition of Assumption 4.2 fails to hold, as the interior of $\mathcal{C}_{\pi}$ is empty, i.e., there exists no $\pi\in\Delta(\mathcal{A})$ such that $V^{g^{j}}(\pi)>\frac{1}{4}$ for all $j\in[4]$. Consequently, there exists no feasible modification $\phi^{i}\in\Phi^{i}_{NM}$ such that $\phi^{i}\circ\pi$ lies in the interior of $\mathcal{C}_{\pi}$. This, in turn, implies that the strong Slater’s condition is also violated.
  2. 2.
    Note that any policy $\pi$ lying on the boundary of $\mathcal{C}_{\pi}$ satisfies $x=y=z=\frac{1}{4}$. For $P1$, all possible deterministic modifications in $\Phi^{i}_{NM,det}$ are given by:  [^3] In the normal-form setting, $\Phi_{NM,det}^{i}=\Phi_{M,det}^{i}$.
    $\phi^{1}(1)(a^{1}\mid 1)=\mathds{1}_{a^{1}=1}\text{ and }\phi^{1}(1)(a^{1}\mid 2)=\mathds{1}_{a^{1}=2}$
    $\phi^{1}(2)(a^{1}\mid 1)=\mathds{1}_{a^{1}=1}\text{ and }\phi^{1}(2)(a^{1}\mid 2)=\mathds{1}_{a^{1}=1}$
    $\phi^{1}(3)(a^{1}\mid 1)=\mathds{1}_{a^{1}=2}\text{ and }\phi^{1}(3)(a^{1}\mid 2)=\mathds{1}_{a^{1}=2}$
    $\phi^{1}(4)(a^{1}\mid 1)=\mathds{1}_{a^{1}=2}\text{ and }\phi^{1}(4)(a^{1}\mid 2)=\mathds{1}_{a^{1}=2}$
    Now, we choose the vector $\alpha=(\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4})$, we have:
    $\sum_{k=1}^{4}\frac{1}{4}V^{g^{j}}(\phi^{i}(k)\circ\pi)=\frac{1}{4},\quad\forall j\in[4].$
    For P2, Assumption 4.2 is satisfied by analogous arguments. Thus, Assumption 4.2 is satisfied.
  3. 3.
    A constrained correlated equilibrium exists since there exists a unique feasible policy, namely $\pi=\left(\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}\right)$. As this is the unique feasible policy, it follows that this policy is a constrained correlated equilibrium.
Appendix C Supporting result for Theorem 4.3
For completeness, we first state Kakutani’s fixed point theorem. Then, we provide a proof of Lemma 5.2 in Section C.1, and a proof of Lemma 5.3 in Section C.2
Theorem C.1 (Kakutani’s Fixed Point [32]).
Let $X$ be a non-empty, compact, and convex subset of a finite-dimensional Euclidean space. Let $F:X\to 2^{X}$ be an upper semi-continuous set-valued function. Then, there exists at least one fixed point $x^{*}$ such that $x^{*}\in F(x^{*})$.
C.1 Proof of Lemma 5.2
In Lemma C.4, we begin by establishing the sensitivity analysis of the linear program $LP^{i}(\pi)$, namely: the upper semi-continuity of its optimal solution set and the continuity of its optimal value with respect to any variability of policy $\pi$ in $LP^{i}(\pi)$. These results will then be used to prove Lemma 5.2.
C.1.1 Sensitivity analysis of $LP^{i}(\pi)$
We recall the linear program $LP^{i}(\pi)$: find $\alpha\in\Delta({K^{i}})$ with $\Phi^{i}_{M,det}=\{\phi^{i}(k)\}_{k=1}^{K^{i}}$ such that:
$\max_{\alpha\in\mathbb{R}^{K^{i}}}\$$\sum_{k=1}^{K^{i}}\alpha_{k}V^{r^{i}}(\phi^{i}(k)\circ\pi)$
s.t. $\sum_{k=1}^{K^{i}}\alpha_{k}V^{g^{j}}(\phi^{i}(k)\circ\pi)\geq c^{j},\,\forall j\in[J],$ (25)
$\sum_{k=1}^{K^{i}}\alpha_{k}=1,$ (26)
$\alpha_{k}\geq 0,\,\forall k\in[K^{i}].$ (27)
For a more compact notation, we recast the above LP such that $\xi(\alpha)\leq 0$ captures the inequality constraints (25) and (27) and $\psi(\alpha)=0$ captures the equality constraints (26). Denote by
$H(\xi,\psi)=\{\alpha\in\mathbb{R}^{K^{i}}\ \lvert\ \xi(\alpha)\leq 0\,\text{ and }\,\psi(\alpha)=0\}$
the set of feasible solutions and by
$M(\pi\mid H(\xi,\psi))=\left\{\alpha\in H(\xi,\psi)\mid\alpha\in\arg\max_{\alpha^{\prime}\in H(\xi,\psi)}\sum_{k=1}^{K^{i}}\alpha_{k}^{\prime}V^{r^{i}}(\phi^{i}(k)\circ\pi)\right\}\subseteq H(\xi,\psi)$
the set of maximizers of the above $LP^{i}(\pi)$.
Next, consider a sequence of policies $\{\pi_{n}\}$ and let $LP_{n}^{i}$, $\xi_{n}$, and $\psi_{n}$ be defined as above, where $\pi$ is replaced by $\pi_{n}$. We make the following assumption which ensures that $LP_{n}^{i}$ and $LP_{n}^{i}$ are feasible and therefore the problem is well-posed.
Assumption C.2.
We assume that both $H(\xi,\psi)$ and $H(\xi_{n},\psi_{n})$ are non-empty.
We furthermore consider the following assumption:
Assumption C.3.
Assume the following holds:
  1. 1.
    For all $j\in[J]$, if there exists a strictly infeasible $\alpha\in\Delta(K^{i})$, i.e.,
    $\sum_{k=1}^{K^{i}}\alpha_{k}V^{g^{j}}(\phi^{i}(k)\circ\pi)<c^{j},$
    then it holds that there does not exist any $\beta\in\mathbb{R}$ such that:
    $\{V^{g^{j}}(\phi^{i}(k)\circ\pi)\}_{k=1}^{K^{i}}=\beta\mathbf{1}_{K^{i}}.$
  2. 2.
    There exists a feasible $\alpha\in\Delta(K^{i})$ with positive weights such that:
    $\sum_{k=1}^{K^{i}}\alpha_{k}V^{g^{j}}(\phi^{i}(k)\circ\pi)\geq c^{j},\,\forall j\in[J],$
    where $\alpha_{k}>0$ for all $k\in[K^{i}]$.
Lemma C.4.
Let Assumptions C.2 and C.3 hold. Assume that $\pi_{n}\to\pi$ pointwise as $n\to\infty$, then the following is satisfied: [^4] Pointwise convergence of $\pi_{n}$ to $\pi$ ensures that for any $i\in\mathcal{N}$, $j\in[J]$, and $k\in[K^{i}]$: $V^{r^{i}}(\phi^{i}(k)\circ\pi_{n})\to V^{r^{i}}(\phi^{i}(k)\circ\pi)\quad\text{and}\quad V^{g^{j}}(\phi^{i}(k)\circ\pi_{n})\to V^{g^{j}}(\phi^{i}(k)\circ\pi)$ pointwise as $n\to\infty$.
  1. 1.
    $\lim\inf_{n\rightarrow\infty}M(\pi_{n}\mid H(\xi_{n},\psi_{n}))\subseteq M(\pi\mid H(\xi,\psi))$.
  2. 2.
    $\lim\sup_{n\rightarrow\infty}M(\pi_{n}\mid H(\xi_{n},\psi_{n}))\subseteq M(\pi\mid H(\xi,\psi))$.
  3. 3.
    The optimal values of $LP_{n}^{i}$ converge to the optimal values of $LP^{i}$.
Proof
  1. 1.
    Denote by $I=\{m\in[J+K^{i}])\ \lvert\ \xi_{m}(\alpha)=0,\,\forall\alpha\in H(\xi,\psi)\}$ the set of those inequality constraints (25) and (27) that are actually equality constraints for all feasible points. Note that this set depends only on the limiting $LP^{i}$. In the following, we will show that:
    ${\lim}\sup_{n\rightarrow\infty}rank(\xi_{n,I},\psi_{n})\leq rank(\xi_{I},\psi).$ (28)
    Suppose for now that Inequality (28) holds, then, by [20, Corollary II.3.4] it follows that either $\lim_{n\rightarrow\infty}H(\xi_{n},\psi_{n})=H(\xi,\psi)$ or $H(\xi_{n},\psi_{n})$ is empty infinitely often. By Assumption C.2 the set $H(\xi_{n},\psi_{n})$ is non-empty for all $n$ and therefore, by [20, Corollary II.3.4] it follows that $\lim_{n\rightarrow\infty}H(\xi_{n},\psi_{n})=H(\xi,\psi)$. Thus, the conditions of [20, Theorem I.2.2] are satisfied and applying it we obtain that $\lim\inf_{n\rightarrow\infty}M(\pi_{n}\mid H(\xi_{n},\psi_{n}))\subseteq M(\pi\mid H(\xi,\psi))$. This concludes the proof of the first point.
    We are now left with showing that Inequality (28) indeed holds.
    Consider $rank(\psi)$ and $rank(\psi_{n})$: Since $\psi_{n}=\psi=\sum_{k=1}^{K^{i}}\alpha_{k}-1$, therefore:
    $rank(\psi)=rank(\psi_{n})=1.$
    Consider $rank(\xi_{I})$ and $rank(\xi_{n,I})$: Consider $\xi:\mathbb{R}^{K^{i}}\rightarrow\mathbb{R}^{J+K^{i}}$ with:
    $\xi_{j}(\alpha)=c^{j}-\sum_{k=1}^{K^{i}}\alpha_{k}V^{g^{j}}(\phi^{i}(k)\circ\pi)\leq 0,\,\forall j\in[J],$
    $\xi_{k}(\alpha)=-\alpha_{k}\leq 0,\,\forall k\in\{J+1,\dots,J+K^{i}\}.$
    Without loss of generality, we assume that $rank(\{\xi_{j}\}_{j\in[J]})=J$ since otherwise the redundant constraints can be removed. Note that removing redundant constraints from $LP^{i}(\pi)$ does not affect $LP^{i}_{n}$ since $I=\{m\in[J+K^{i}]\ \lvert\ \xi_{m}(\alpha)=0,\,\forall\alpha\in H(\xi,\psi)\}$ depends only on the limiting $LP^{i}(\pi)$. Next, we divide the satisfaction of each constraint $\xi_{j}$ for $j\in[J]$ into three distinct cases:
    1. (a)
      For all $\alpha\in\Delta(K^{i})$, we have $\xi_{j}(\alpha)=0$.
    2. (b)
      For all $\alpha\in\Delta(K^{i})$, we have $\xi_{j}(\alpha)\geq 0$. Furthermore, there exists a feasible $\alpha\in\Delta(K^{i})$ and a strictly infeasible solution $\alpha^{\prime}\in\Delta(K^{i})$, i.e., $\xi_{j}(\alpha)=0$ and $\xi_{j}(\alpha^{\prime})>0$, respectively.
    3. (c)
      There exists a strictly feasible solution $\alpha\in\Delta(K^{i})$ such that $\xi_{j}(\alpha)<0$, hence $j\notin I$.
    In case (a), the constraint can be removed from the linear program $LP^{i}(\pi)$ as it does not affect the feasible set. Without loss of generality, assume that there are $J_{1}\in[J]$ constraints $\{\xi_{j}\}$ for $j\in[J_{1}]$ satisfying condition (b), and the remaining $J-J_{1}$ constraints $\{\xi_{j}\}$ for $j=J_{1}+1,\dots,J$ satisfy condition (c). Therefore, we have $J_{1}]\subseteq I$.
    For the additional inequality constraints $\xi_{k}$, where $k=J+1,\dots,J+K^{i}$, Assumption C.3 guarantees the existence of a feasible $\alpha\in\Delta(K^{i})$ with positive weights, implying again that $I=[J_{1}]$.
    Suppose that $(\xi_{I},\psi)$ has full rank, i.e., $\mathrm{rank}(\xi_{I},\psi)=J_{1}+1$. Then, the following inequality holds:
    $\mathrm{rank}(\xi_{n,I},\psi_{n})\leq\mathrm{rank}(\xi_{I},\psi),\quad\forall n,$
    since $\mathrm{rank}(\xi_{n,I},\psi_{n})\leq J+1$. Hence, Inequality (28) is satisfied.
    Now, we show that $\mathrm{rank}(\xi_{I},\psi)=J_{1}+1$. Any linear mapping $(\xi_{I},\psi):\mathbb{R}^{K^{i}}\to\mathbb{R}^{J_{1}+1}$ can be represented as $\alpha^{\top}M-(1,c)^{\top}$, where $M\in\mathbb{R}^{(J_{1}+1)\times K^{i}}$ is defined as:
    $M:=\begin{bmatrix}1&1&\dots&1\\ V^{g^{1}}(\phi^{i}(1)\circ\pi)&V^{g^{1}}(\phi^{i}(2)\circ\pi)&\dots&V^{g^{1}}(\phi^{i}(K^{i})\circ\pi)\\ \vdots&\vdots&\ddots&\vdots\\ V^{g^{J_{1}}}(\phi^{i}(1)\circ\pi)&V^{g^{J_{1}}}(\phi^{i}(2)\circ\pi)&\dots&V^{g^{J_{1}}}(\phi^{i}(K^{i})\circ\pi)\end{bmatrix}$.
    By Assumption C.3, we have $\mathrm{rank}(M)=J_{1}+1$, which concludes the proof.
  2. 2.
    The proof follows from the first point and the same arguments as in [2, Theorem 3.6 (ii)].
  3. 3.
    The proof follows from the first point and the same arguments as in [2, Theorem 3.6 (iv)].
C.1.2 Main proof of Lemma 5.2
We are ready to provide a proof of Lemma 5.2.
Proof We will prove the properties one by one:
  1. 1.
    Convexity and compactness are standard results of linear programming [12]. Upper semi-continuity of $\Omega^{i}(\pi)$ follows from Lemma C.4, if we can verify that the conditions of the lemma are satisfied:
    Assumption C.2:
    This holds since for any feasible policy $\pi$ the identity is a feasible modification.
    Assumption C.3:
    We need to verify that Assumption C.3 holds for all $\pi\in\mathcal{C}_{\pi}$ and all $i\in\mathcal{N}$.
    For any policy lying in the interior of $\mathcal{C}_{\pi}$, the identity modification $\phi^{i}(\mathrm{id})\in\Phi^{i}_{\mathrm{M,det}}$ always ensures $V^{g^{j}}(\phi^{i}(\mathrm{id})\circ\pi)>c^{j}$. Now, pick any $\alpha^{\prime}\in\Delta(K^{i})$ such that $\alpha^{\prime}_{k}>0$ for all $k\in[K^{i}]$, and define $\alpha^{\prime\prime}:=(1-\varepsilon)\mathbf{1}_{\mathrm{id}}+\varepsilon\alpha^{\prime}$, where $\mathbf{1}_{\mathrm{id}}$ is the canonical basis vector whose position of $1$ corresponds to the identity modification. By choosing $\varepsilon>0$ sufficiently small, it follows that $\sum_{k=1}^{K^{i}}\alpha^{\prime\prime}_{k}V^{g^{j}}(\phi^{i}(k)\circ\pi)>c^{j}$ with $\alpha^{\prime\prime}_{k}>0$ for all $k\in[K^{i}]$.
    For policies on the boundary of $\mathcal{C}_{\pi}$, if there exists a strictly infeasible $\alpha\in\Delta(K^{i})$ such that
    $\sum_{k=1}^{K^{i}}\alpha_{k}V^{g^{j}}(\phi^{i}(k)\circ\pi)<c^{j}.$
    Then, there must exist at least one modification $\phi^{i}(k)$ for which $V^{g^{j}}(\phi^{i}(k)\circ\pi)<c^{j}$. On the other hand, under the identity modification $\phi^{i}(\mathrm{id})$, we have $V^{g^{j}}(\phi^{i}(\mathrm{id})\circ\pi)=c^{j}$. Thus, there does not exist any $\beta\in\mathbb{R}$ such that $\{V^{g^{j}}(\phi^{i}(k)\circ\pi)\}_{k=1}^{K^{i}}=\beta\mathbf{1}_{K^{i}}$. Hence, together with Assumption 4.2, Assumption C.3 is satisfied.
C.2 Proof of Lemma 5.3
In the following, we provide a proof of Lemma 5.3.
Proof Consider any element $\{d^{i}_{t}\}_{t\in[H]}\in A^{i}(\pi)$, defined as:
$d_{t}^{i}(s,a):=\sum_{k=1}^{K^{i}}\alpha_{k}d_{t}^{\phi^{i}(k)\circ\pi}(s,a),\,\forall(s,a)\in\mathcal{S}\times\mathcal{A},\,\forall t\in[H],$ (29)
where $\alpha\in\Omega^{i}(\pi)$. For all $\pi^{\prime}\in\Gamma(d^{i})$, we have that:
$V^{g^{j}}(\pi^{\prime})=$ $\sum_{t=1}^{H}\sum_{s_{t},a_{t}}d_{t}^{i}(s_{t},a_{t})g_{t}^{j}(s_{t},a_{t})$
$\overset{(i)}{=}$ $\sum_{t=1}^{H}\sum_{s_{t},a_{t}}\sum_{k=1}^{K^{i}}\alpha_{k}d_{t}^{\phi^{i}(k)\circ\pi}(s_{t},a_{t})g_{t}^{j}(s_{t},a_{t})$
$=$ $\sum_{k=1}^{K^{i}}\alpha_{k}\sum_{t=1}^{H}\sum_{s_{t},a_{t}}d_{t}^{\phi^{i}(k)\circ\pi}(s_{t},a_{t})g_{t}^{j}(s_{t},a_{t})$
$=$ $\sum_{k=1}^{K^{i}}\alpha_{k}V^{g^{j}}(\phi^{i}(k)\circ\pi)$
$\overset{(ii)}{\geq}$ $c^{j},$
where $(i)$ uses Equation (29) and $(ii)$ holds since $\alpha$ is a solution of $\Omega^{i}(\pi)$. Thus, $d^{i}\in\mathcal{C}_{d}$, implying that $A^{i}(\pi):\mathcal{C}_{\pi}\to 2^{\mathcal{C}_{d}}$. Note that $\sum_{k=1}^{K^{i}}\alpha_{k}d_{t}^{\phi^{i}(k)\circ\pi}(s,a)$ is polynomial in $\pi$ for all $t\in[H]$ and $(s,a)\in\mathcal{S}\times\mathcal{A}$ and thus continuous in $\pi$. Furthermore, Lemma 5.2 together with the fact that $A^{i}(\pi)$ is a linear transformation of $\Omega^{i}(\pi)$ with respect to $\pi$ ensures that $A^{i}(\pi)$ maps to a convex and compact set in $2^{\mathcal{C}_{d}}$, and is upper semi-continuous for any $\pi\in\mathcal{C}_{\pi}$.